### v1

# Roundtable Policy: Improving Scientific Reasoning and Narratives through Confidence-Weighted Consensus of LLMs

Abstract
Large language models (LLMs) have demonstrated remarkable capabilities not only in language generation but also in advancing scientific discovery. A growing body of work has explored ways to improve their reasoning, from self-consistency and chain-of-thought to multi-agent debate. Inspired by the dynamics of scientific committees and The Society of Mind, we introduce Roundtable Policy, a complementary inference-time reasoning framework that performs inference through the weighted consensus of multiple LLMs. Our findings indicate that this approach significantly enhances reasoning in complex heterogeneous scientific tasks and improves scientific narratives in terms of creativity, rigor, and logical coherence, while reducing hallucinations that single models are prone to. Our approach emphasizes structured and interpretable consensus rather than opaque convergence, while requires only black-box access and uniform procedures, making it broadly applicable to multi-LLM reasoning. Project website at https://ai4scidisc.github.io/roundtable/.
1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities in text generation, reasoning, and few-shot adaptation. These advances have opened exciting opportunities for scientific discovery, where complex, interdisciplinary problems demand not only accurate reasoning but also coherent narrative construction. However, current LLMs may hallucinate facts, introduce implausible logical jumps, or produce fragmented explanations. Such limitations are particularly problematic in scientific contexts, where reliability and clarity are equally critical. For instance, a scientific proposal must not only provide factually correct content, but also organize it into a structured and persuasive narrative. Addressing these dual requirements calls for mechanisms that go beyond the capacity of a single LLM.
A natural motivation is therefore to leverage multiple LLMs, aggregating their complementary strengths. Prior work has explored self-consistency with multiple sampled reasoning paths (Wang et al., 2022), multi-agent debate frameworks (Du & et al., 2023; Madaan & et al., 2023), and tool-augmented multi-agent systems (Schick et al., 2023; Ghafarollahi & Buehler, 2024). However, these approaches are often limited to single-task reasoning, rely on emergent dynamics that are difficult to interpret, or lack systematic mechanisms for weighting contributions across heterogeneous domains. Ensemble-style methods such as majority voting treat all models equally, overlooking that different models excel in different domains or evaluation dimensions. Debate-style methods, while sometimes converging to better outputs, remain inherently opaque: it is often unclear why a particular answer was selected, or hard to quantify which models were more trustworthy in the process. In scientific applications, this lack of transparency poses a major barrier, since researchers require not only correct outputs but also interpretable evidence supporting the decision.
At a broader level, scientific problem solving is inherently multi-faceted, requiring both precise reasoning and coherent long-form narratives. While individual LLMs have made impressive progress, they still face challenges in consistently meeting these diverse demands across domains and tasks. This motivates the need for a structured approach to harness collective intelligence—one that integrates complementary expertise, surfaces and resolves disagreements transparently, and preserves accountability in the decision process.
Building on these considerations, we introduce Roundtable Policy, an inference-time framework that formalizes deliberation among multiple LLMs into a principled process for trustworthy scientific reasoning and narrative generation (Figure 1). In our approach, diverse models first propose candidate responses, which are then jointly evaluated to record their performances. These evaluations are distilled into a confidence-weight table that encodes both historical reliability and uncertainty for each candidate across tasks and dimensions. At inference, this confidence-weight table guides a fusion process that integrates the diverse responses into a single consensus answer—balancing factual accuracy with narrative coherence while preserving transparency in how the decision is reached. As illustrated in Figure 1, even on a scientific question such as explaining why nucleation accelerates at greater undercooling while growth dominates at smaller undercooling, individual models provide partial or inconsistent accounts. The roundtable process fuses these perspectives into a coherent explanation, showing how consensus-driven reasoning can elevate fragmented insights into a more complete scientific narrative.
We evaluate Roundtable Policy in two complementary scientific regimes. In the MultiTask setting, heterogeneous problems from geoscience, biology, mathematics, and materials science are solved simultaneously, testing cross-domain reasoning. In the SingleTask setting, models generate structured scientific proposals, testing long-form narrative construction. Across both, our findings show that Roundtable Policy consistently improves factual reliability, reasoning quality, and narrative coherence, yielding an average 13.01% improvement in MultiTask and 11.04% improvement in SingleTask compared to individual models.
Unlike ensemble-style methods such as majority voting, which treat all models equally, our approach adaptively weights models based on demonstrated reliability. And in contrast to debate-style frameworks, whose convergence dynamics are opaque, Roundtable Policy produces explicit, auditable signals that make the aggregation process transparent. This combination of principled weighting and interpretability distinguishes our method as a scalable and accountable alternative for multi-model scientific reasoning.
In summary, our contributions are threefold: (i) we propose Roundtable Policy, an inference-time framework for performance-weighted consensus aggregation among multiple LLMs; (ii) we demonstrate that this approach enhances both factual reasoning and scientific narrative across heterogeneous tasks; and (iii) we release a graduate-level materials science benchmark derived from UCLA PhD qualifying exams, which serves as a rigorous, domain-authentic testbed and complements existing evaluations of scientific LLMs. More broadly, the principles underlying Roundtable Policy apply across multi-model scientific workflows, offering an interpretable and generalizable path toward AI-assisted reasoning and communication in science.
2 Methodology
This section formalizes the Roundtable Policy framework. We begin with the intuition behind the Roundtable Policy and then present the task settings and notation (a complete list of symbols is provided in Appendix §A). We next describe how the confidence-weight table is learned from repeated interactions between players and graders, and how it subsequently guides inference through consensus-weighted fusion.
2.1 Design Intuitions behind Roundtable Policy
Scientific decision-making rarely relies on a single voice. In research committees, for example, each member independently evaluates a proposal, but the final judgment is not determined by a simple vote. Instead, reviewers’ assessments are weighed by domain expertise, reliability, and the clarity of justification before being synthesized into a collective decision. This process balances agreement with accountability, ensuring that the outcome is both credible and transparent.
Roundtable Policy draws on the same principle. Rather than treating LLMs as isolated predictors, we view them as members of a committee. A set of players propose candidate responses informed by diverse inductive biases. These are evaluated by an ensemble of graders, whose rubric-based assessments provide explicit signals of quality and uncertainty. Finally, a structured fusion step aggregates these evaluations into a single consensus output—analogous to how a committee chair synthesizes diverse reviews into a reasoned decision.
This design is further motivated by the perspective of The Society of Mind (Minsky, 1988), which frames intelligence as the coordination of many specialized, relatively simple agents. By embracing this philosophy, Roundtable Policy transforms the inherent diversity of LLMs into a strength, enabling collective reasoning that is both interpretable and robust.
2.2 Task Settings
We evaluate Roundtable Policy in two complementary scientific regimes that mirror the dual demands of scientific reasoning: solving pointwise problems and constructing coherent narratives.
MultiTask. Models are simultaneously queried with nine heterogeneous scientific questions spanning geoscience, biology, mathematics, and materials science (with six distinct sub-categories). Task formats include multiple-choice, fill-in-the-blank, and short-answer questions. Intuitively, this setting resembles a written exam with diverse problem types—success requires breadth of expertise rather than mastery of a single domain.
SingleTask. Models are prompted to generate structured scientific proposals following a standardized three-paragraph template with sections for Background, Methodology, and Impact. This setting captures the narrative aspect of scientific work: beyond producing isolated facts, one must weave them into a logically consistent and persuasive storyline.
Across both regimes, we deploy pretrained LLMs (players) to generate candidate responses and LLMs (graders) to form the AI Committee. Just as in a real committee, players provide diverse answers while graders act as evaluators who judge each contribution against shared standards. In MultiTask, graders jointly evaluate the responses per round, analogous to examiners scoring a large pool of student answers. In SingleTask, graders assess proposals paragraph-by-paragraph along three rubric dimensions—creativity, scientific rigor, and logical coherence—reflecting the multifaceted criteria by which scientific proposals are judged.
Each evaluation yields a trust weight with two components: (i) a quality score in , where negative values indicate incorrect or misleading responses, and (ii) a calibrated 95% confidence interval representing grader certainty. Together, these form the equivalent of both a “grade” and a measure of “grading confidence,” ensuring that downstream decisions reflect not only evaluation outcomes but also their reliability. Notation follows Appendix §A denotes the -th player agent, the -th grader, and the fusion agent. Player responses and grader scores are indexed by round .
2.3 Learning the Confidence-Weight Table
The confidence-weight table serves as a structured memory of each player’s historical reliability across subtasks (MultiTask) or rubric dimensions (SingleTask). Unlike fine-tuning approaches, Roundtable Policy does not alter the parameters of the underlying LLMs. Instead, is iteratively updated from repeated rounds of evaluation, much like how a committee gradually learns which members are most reliable in which topics.
At round , a query is sampled from the task dataset. Players generate responses , which are pooled into . Graders then evaluate these responses, producing , which are averaged across graders to yield . After rounds, the cumulative evaluations are distilled into:
capturing both the estimated quality and the associated uncertainty of each player. Intuitively, this table plays the role of a “track record” or “reputation system” for each player, ensuring that strong but inconsistent voices are weighted differently from consistently reliable ones. This design calibrates the committee’s trust in different players without retraining them, enabling adaptability and efficiency.
2.4 Inference with the Confidence-Weight Table
Once has been learned, it guides consensus fusion at inference. Given a new query , each player generates a response . The collection is passed to the fusion process , which aggregates them using the pretrained :
Here, ensures that expertise is applied where it matters most. In MultiTask, it emphasizes domain-specific reliability (e.g., assigning higher weights to players consistently strong in mathematics). In SingleTask, it captures stylistic and analytical strengths across rubric dimensions, reflecting how different voices contribute to different aspects of a proposal.
The fusion process itself is not fine-tuned; it acts as an orchestrator that conditions on to produce a consensus answer. This mimics how a committee chair integrates both the content and credibility of reviewers before issuing a final recommendation. Optionally, the AI Committee can re-evaluate to assign a quality score and confidence interval, preserving interpretability and accountability throughout the process.
3 Experiments
We design experiments to systematically evaluate the effectiveness and robustness of the proposed Roundtable Policy. Our analysis is guided by the following research questions:
RQ1: To what extent does Roundtable Policy enhance scientific reasoning and writing performance relative to individual LLMs? (see §3.1)
RQ2: How stable and robust is Roundtable Policy over long-run evaluation regimes, and does it consistently provide more reliable performance? (see §3.2)
RQ3: To what degree do AI graders introduce bias, and can the AI Committee mitigate these biases to achieve greater inter-grader consistency? (see §3.3)
3.1 Performance Improvement (RQ1)
| Setting | Model | Material Science Dataset | Others | |||||||
| Mech. 1 | Mech. 2 | EE | Opt. | Thermo. | Semi. | Math | Geo. | Bio. | ||
MultiTask |
GPT-4o | 85.894.27 | 78.265.37 | 87.714.11 | 74.595.49 | 85.144.40 | 85.584.60 | 87.812.13 | 83.552.20 | 51.536.64 |
| Grok-2 | 83.924.07 | 76.244.96 | 85.613.92 | 76.184.99 | 82.024.10 | 82.634.39 | 88.841.70 | 77.392.28 | 51.316.25 | |
| Claude-3.5H | 82.074.65 | 71.206.04 | 84.724.37 | 71.495.83 | 77.745.03 | 77.955.22 | 81.192.58 | 73.682.58 | 48.526.75 | |
| GPT-4.5p | 72.866.77 | 71.466.96 | 76.676.12 | 72.786.36 | 77.185.71 | 77.686.17 | 82.502.68 | 82.592.08 | 50.636.89 | |
| DeepSeek-chat | 77.635.35 | 72.026.14 | 81.534.97 | 72.965.79 | 79.644.87 | 79.815.23 | 72.233.16 | 78.052.41 | 47.836.57 | |
| Claude-3.7S | 77.433.46 | 70.514.05 | 79.243.13 | 72.393.70 | 77.413.49 | 79.493.12 | 62.342.93 | 72.561.97 | 47.765.36 | |
| Llama-3.3 | 72.145.96 | 63.866.95 | 75.335.55 | 67.136.30 | 71.235.71 | 73.875.72 | 87.431.93 | 72.772.49 | 48.646.54 | |
| Qwen-Plus | 80.325.01 | 72.366.03 | 82.974.71 | 73.425.66 | 81.494.70 | 81.184.93 | 86.992.08 | 76.272.48 | 51.646.57 | |
| Gemini-2.0 | 67.267.61 | 61.618.09 | 74.966.58 | 65.737.31 | 70.726.68 | 71.606.99 | 49.774.79 | 75.172.58 | 42.117.31 | |
| RP | 87.644.26 | 81.505.36 | 89.994.00 | 81.294.96 | 86.554.09 | 91.824.34 | 94.801.01 | 85.331.63 | 47.557.02 | |
| Setting | Model | Background | Methodology | Impact | ||||||
| Crea. | Sci. Rig. | Log. Coh. | Crea. | Sci. Rig. | Log. Coh. | Crea. | Sci. Rig. | Log. Coh. | ||
SingleTask |
GPT-4o | 70.995.01 | 74.734.68 | 82.753.76 | 75.495.23 | 74.815.03 | 79.634.31 | 72.155.19 | 70.535.25 | 79.644.19 |
| GPT-o3 | 84.073.98 | 84.253.82 | 86.463.42 | 85.923.95 | 86.703.56 | 85.843.56 | 83.904.15 | 81.524.25 | 84.863.69 | |
| GPT-4.1 | 76.964.62 | 81.374.16 | 86.193.51 | 81.654.67 | 83.194.19 | 84.753.85 | 78.814.74 | 77.914.67 | 83.903.88 | |
| GPT-4.1n | 77.394.55 | 80.944.15 | 85.593.53 | 80.484.69 | 82.304.19 | 83.973.85 | 77.974.74 | 77.224.65 | 82.943.90 | |
| GPT-o3-m | 82.654.16 | 84.783.83 | 87.263.36 | 83.154.35 | 86.863.53 | 86.443.53 | 81.504.40 | 81.824.23 | 85.153.67 | |
| GPT-4o-m | 72.784.94 | 74.644.70 | 82.193.80 | 75.445.16 | 76.184.85 | 79.814.23 | 72.785.18 | 70.695.24 | 78.824.25 | |
| Grok-3 | 78.505.10 | 79.104.50 | 85.503.64 | 80.964.75 | 81.264.47 | 83.583.99 | 78.814.78 | 76.344.93 | 83.044.01 | |
| Claude-3.7S | 82.474.22 | 84.733.92 | 88.173.28 | 85.524.20 | 86.653.79 | 86.893.57 | 84.534.22 | 82.134.38 | 86.413.59 | |
| DeepSeek-chat | 84.054.20 | 83.744.12 | 88.283.19 | 84.214.31 | 84.584.13 | 86.083.58 | 83.684.23 | 81.564.48 | 86.503.46 | |
| RP | 90.253.95 | 89.953.78 | 91.043.09 | 90.393.91 | 91.283.35 | 91.183.20 | 90.153.95 | 89.863.95 | 90.993.26 | |
To answer RQ1, we compare Roundtable Policy (RT) against individual LLM baselines across both MultiTask and SingleTask settings. Table 1 summarizes overall performance, while Figures 2(a)–2(b) visualize relative gains. This correspondence between player ID and model follows Table 3. The correspondence of subtasks of MultiTask and task IDs follows Table 4.
Quantitative Analysis.
In the MultiTask regime, RT delivers a 13.01% average lift, outperforming 8 out of 9 single-model baselines. The only exception is Task 9 (Bio.), where Qwen-plus () and GPT-4o () slightly exceed RT (). Beyond accuracy, RT reduces uncertainty; for instance, on Task 7 (Math) the confidence interval contracts to compared with baselines’ –. Figure 2(a) further shows per-player improvements. Gains are broadly positive on Tasks 1–8, with especially large margins on reasoning-intensive subtasks—for example, +90.5% over Gemini-2.0 on Task 7. Smaller improvements occur when a single model already aligns closely with ground truth, while Task 9 shows marginal or negative differences—consistent with its per-task averages. Appendix §D.1 confirms that weaker players (e.g., Gemini-2.0) enjoy the largest lifts, while stronger models exhibit limited headroom.
In the SingleTask regime, which evaluates structured scientific proposal writing, RT achieves the best performance in all nine rubric cells, explaining the observed 11.04% average lift. Each proposal comprises three sections—Background, Methodology, and Impact—scored across Creativity, Scientific Rigor, and Logical Coherence. Scores cluster tightly around 90–91% with narrow confidence intervals (typically –). Gains concentrate in the Impact section, particularly Scientific Rigor, as shown in Figure 2(b). Appendix §D.2 reveals positive medians across most players, with right-skewed tails indicating that even when consensus does not outperform the strongest model, it still elevates weaker drafts significantly.
Qualitative Analysis.
These results illustrate how RT acts like a committee process rather than a flat average. In MultiTask, the confidence-weight table down-weights unreliable contributions and amplifies complementary strengths, thereby “rescuing” weaker players without diluting strong ones. This mirrors how, in a committee, flawed proposals can still contribute useful fragments when evaluated critically. The large gains on reasoning-heavy tasks (e.g., Tasks 6–7) highlight how diverse error patterns across models can be combined into substantially stronger reasoning chains.
In SingleTask writing, weaker players show the most pronounced gains in Impact and Methodology. For instance, GPT-4o exhibits long right tails in Impact–Scientific Rigor and Impact–Logical Coherence, while GPT-o3 shows consistently positive medians with bounded variance. Even strong baselines such as DS-chat rarely degrade under RT, suggesting that consensus seldom harms high-quality drafts but frequently elevates weaker ones. Mechanistically, the confidence-weight table enforces rubric-aware, non-redundant fusion: systematically weak contributors are down-weighted, while demonstrated strengths are amplified. The pattern resembles how peer reviewers specialize—some focusing on methodology, others on impact—and the editor synthesizes these perspectives into a well-rounded final decision.
3.2 Long-Run Stability and Robustness (RQ2)
Setting Player ID LLM # Mean Gain (%) Std (%) p-value Significance MultiTask 001 GPT-4o 36000 6.476865 77.280369 ✓ 002 Grok-2 36000 5.254293 78.231667 ✓ 003 Claude-3.5Haiku 36000 17.251754 101.839946 ✓ 004 GPT-4.5-preview 36000 18.552776 94.463211 ✓ 005 Deepseek-Chat 36000 16.048551 96.935292 ✓ 006 Claude-3.7Sonnet 36000 13.732040 121.321577 ✓ 007 Llama-3.3 36000 24.988341 123.678543 ✓ 008 Qwen-plus 36000 12.586367 91.751670 ✓ 009 Gemini-2.0 36000 46.474125 145.071773 ✓ SingleTask 001 GPT-4o 36000 27.087781 40.303119 ✓ 002 GPT-o3 36000 13.316247 148.785675 ✓ 003 GPT-4.1 36000 15.107889 73.600957 ✓ 004 GPT-4.1-nano 36000 15.805949 28.188351 ✓ 005 GPT-o3-mini 36000 9.687830 21.760829 ✓ 006 GPT-4o-mini 36000 27.206693 42.707340 ✓ 007 Grok-3 36000 16.485277 29.024328 ✓ 008 Claude-3.7Sonnet 36000 7.761157 16.993190 ✓ 009 Deepseek-Chat 36000 7.165863 7.474061 ✓
To answer RQ2, we test whether RT’s gains remain statistically reliable over long-run evaluations. Using the Wilcoxon signed-rank test (§A.2) on round-level paired differences, RT consistently outperforms all baselines. As shown in Table 2, across all 18 models RT achieves significant improvements at , with -values typically below .
Quantitative Analysis.
In the MultiTask regime, RT consistently outperforms baselines with mean percentage gains ranging from +5.3% (Grok-2) to +46.5% (Gemini-2.0). Even strong players such as GPT-4o (+6.5%) and Claude-3.7Sonnet (+13.7%) see reliable gains, highlighting that consensus remains beneficial even when individual agents are already competitive. The large improvement over Gemini-2.0 illustrates how RT effectively leverages complementary strengths to elevate weaker contributors, rather than letting them dilute overall performance.
In the SingleTask regime, RT delivers significant lifts across all nine baselines, with mean improvements from +7.2% (Deepseek-Chat) to +27.2% (GPT-4o-mini). Notably, gains are especially pronounced for smaller or lighter-weight models (e.g., GPT-4o-mini, GPT-4o), suggesting that RT acts as a stabilizing mechanism that brings their drafts closer to the quality of stronger models. Even against advanced systems like GPT-o3 and Claude-3.7Sonnet, improvements remain consistently positive, confirming robustness across a wide range of capability levels.
Qualitative Analysis.
Beyond statistical significance, the long-run results highlight important qualitative patterns. In the MultiTask regime, RT’s robustness stems from its ability to repeatedly exploit complementary error profiles: weaker players such as Gemini-2.0 or Llama-3.3 often fail in different ways, but when their contributions are weighted through the confidence table, the ensemble consistently extracts correct elements while suppressing noise. This is reflected in the large variance reported in Table 2—although individual rounds fluctuate, the aggregated effect over thousands of rounds converges to stable gains. In committee terms, even if certain members are unreliable in a given discussion, their occasional correct insights are preserved while systematic mistakes are filtered out, ensuring dependable group performance.
In the SingleTask regime, stability is most visible at the rubric level. Analytical dimensions such as Impact–Scientific Rigor show the largest sustained gains, whereas stylistic traits (e.g., Creativity or Logical Coherence) improve more modestly. This asymmetry suggests that RT is especially effective in consolidating rigorous, reasoning-heavy contributions—akin to how editors place greater weight on technical soundness than on stylistic flair when synthesizing reviewer reports. Importantly, the consistency of improvements across tens of thousands of rounds demonstrates that RT does not rely on one-off synergies, but instead behaves like a well-tuned review board: repeatedly integrating diverse perspectives into a balanced and accountable consensus.
3.3 Bias and Consistency of AI Graders (RQ3)
To address RQ3, we examine grader bias and inter-grader consistency through two complementary lenses: (i) the distribution of ranks assigned to RT by different graders, and (ii) pairwise agreement measured by Kendall’s Tau (see §A.3). Aggregated results are shown in Figure 3, with task- and dimension-level breakdowns provided in Figures 14 and 13 in §D.3.
Quantitative Analysis.
Figure 3 highlights a clear distinction between MultiTask and SingleTask regimes. In MultiTask, the aggregated rank distribution is steep and concentrated in the top three positions, with RT frequently ranked first. Correspondingly, Kendall’s Tau values are consistently high (–), reflecting strong alignment among graders. Task-level inspection (Figures 14, 13) confirms this pattern: Tasks 001–006 exhibit overlapping distributions, Tasks 007–008 achieve near-perfect consensus (), while Task 009 shows the weakest agreement (–) as some graders disperse RT’s rank across mid-range positions.
In SingleTask, however, the aggregated rank distribution is flatter and more dispersed, and Kendall’s Tau values drop markedly (– on average). Creativity-related dimensions (e.g., Background–Creativity, Methodology–Creativity) place RT near the top but with substantial grader variability. Logical coherence dimensions show the lowest alignment, with Tau values as low as –, indicating inconsistent judgments. By contrast, Impact–Scientific Rigor yields the strongest agreement, with RT concentrated near rank 1 and Tau values of –. Together, these results demonstrate that graders achieve high consensus in objective, structured tasks but diverge significantly when assessing subjective, open-ended proposals.
Qualitative Analysis.
These results reveal that bias is minimal in structured MultiTask settings, where evaluation criteria are well defined, but more pronounced in open-ended SingleTask writing, where subjective interpretation shapes grading. This mirrors human peer review: reviewers evaluating technical correctness tend to agree, while judgments about creativity or style diverge. Crucially, the AI Committee mitigates such variability by aggregating across graders, reducing the impact of individual biases. In practice, this means RT functions like an editorial board: even when reviewers disagree, their collective judgment yields a more balanced and reliable outcome.
4 Ablation Study
To disentangle the contributions of our design choices and validate their scientific relevance, we conduct two ablation studies. These analyses illustrate how Roundtable Policy strengthens structured reasoning and improves narrative quality by highlighting what is gained—and what is lost—when key components are removed.
Policy Ablation.
We first compare RT with a Best-Single-Model Policy (BSMP), which simply selects the response from the single model with the highest confidence weight rather than synthesizing across multiple players. Conceptually, BSMP resembles a committee that crowns one “top performer” and ignores all other voices. The results in Figure 4 make clear why this approach falls short: across both MultiTask and SingleTask settings, RT consistently outperforms BSMP. In MultiTask, the advantages are particularly evident in reasoning-heavy domains such as mathematics and mechanics, where drawing on complementary perspectives yields large gains. In SingleTask proposal writing, consensus fusion not only raises average scores across all rubric dimensions but also reduces variance, producing outputs that are both stronger and more consistent. These findings highlight that scientific progress rarely emerges from a single dominant voice; it is the synthesis of multiple viewpoints that produces reliable reasoning and coherent narrative.
Rounds Ablation.
We next examine the impact of the number of training rounds used to construct the confidence-weight table. Here, the table plays the role of institutional memory—accumulating evidence about which players are dependable under different conditions. As shown in Figure 5, the effect of this memory differs by task type. In MultiTask, where benchmarks span diverse scientific domains, performance steadily improves as more rounds are incorporated and stabilizes only after hundreds of evaluations. By contrast, SingleTask writing converges much more quickly: fewer than 100 rounds already suffice to surpass baselines, with further rounds offering only modest additional benefit. This divergence reflects a natural distinction: heterogeneous domains require extensive calibration to build trust across disciplines, whereas rubric-based writing tasks benefit from faster stabilization since evaluative criteria are more homogeneous.
Summary.
These ablations underscore two core insights: First, weighted consensus is fundamentally more powerful than relying on a single “best” model, as robust science arises from collective deliberation rather than solitary voices. Second, the effectiveness of the confidence-weight table depends on task structure: structured MultiTask settings demand long-run calibration, while open-ended SingleTask writing achieves stability with much less historical data. Together, these findings reinforce that consensus-driven fusion is not only effective but also principled—mirroring how real scientific communities build reliable knowledge through collective judgment.
5 Related Work
LLMs for Scientific Discovery. Large language models (LLMs) have rapidly expanded the scope of data-driven science, from AlphaFold’s breakthroughs in structural biology (Jumper et al., 2021) to recent applications across mathematics (Yue et al., 2024), chemistry (Yu et al., 2024), biology (Labrak et al., 2024), and geography (Li et al., 2023). Their strength lies in structured reasoning and seamless integration with external tools, enabling automated workflows (Cao, 2017). Yet, complex multi-step reasoning often requires specialized prompting, tool-augmented inference, or fine-tuning, motivating the rise of agentic systems that combine reasoning with external solvers and databases (Boiko et al., 2023; Lu et al., 2024).
Reasoning and Factuality in Language Models. Improving reasoning and factuality in LLMs has been studied extensively. Reasoning can be enhanced through prompting methods such as scratchpads (Nye & et al., 2021), chain-of-thought demonstrations (Wei et al., 2022; Kojima et al., 2022; Reynolds & McDonell, 2021), verification and self-consistency (Cobbe et al., 2021; Wang et al., 2022), and reflection (Shinn et al., 2023; Madaan & et al., 2023), as well as fine-tuning approaches (Rajani et al., 2019; Zelikman et al., 2022; Lewkowycz et al., 2022). Factuality, in turn, has been strengthened through RLHF (Christiano et al., 2017; Ziegler & et al., 2019; Ouyang et al., 2022), dataset curation (Lee & et al., 2022), retrieval-based augmentation (Guu et al., 2020), and likelihood-based calibration (Kadavath & et al., 2022).
Multi-Agent and Ensemble Approaches. Beyond single-model prompting, researchers have explored ensemble and multi-agent strategies to enhance reliability. Early work shows that self-consistency (Wang et al., 2022) and majority voting (Li et al., 2022; Cobbe et al., 2021; Thoppilan et al., 2022) can mitigate errors by aggregating diverse reasoning chains. Building on this, multi-agent debate frameworks (Irving et al., 2018; Du & et al., 2023; Madaan & et al., 2023) and collaborative dialogue methods (Chan & et al., 2023; Wang & et al., 2023) let models iteratively critique and refine one another’s outputs. Related lines of work also investigate compositional generation, combining outputs from multiple models (Du et al., 2020; Liu & et al., 2022; Alayrac & et al., 2022; Zeng & et al., 2022). While these methods demonstrate the value of collective intelligence, they often rely on flat aggregation or emergent convergence. In contrast, our approach introduces a structured consensus mechanism that explicitly weights models by demonstrated reliability, producing interpretable and auditable signals of trust.
Benchmarks for LLM-Based Agents. Evaluating agentic LLMs has inspired benchmarks ranging from synthetic tool-use tasks (e.g., TaskBench (Shen et al., 2024)) to domain-specific challenges such as SWE-Bench (Jimenez et al., 2024), BioCoder-Py (Tang et al., 2024), and MLAgentBench (Huang et al., 2024). More recently, datasets sourced from scientific literature (Majumder et al., 2024) aim to capture authentic workflows. Our work complements these efforts by introducing a domain-authentic benchmark in materials science that stresses both mechanistic understanding and property prediction, thereby testing agents under both interpretative and predictive dimensions.
Machine Learning and Agents for Materials Science. Machine learning has already transformed materials science by enabling scalable property prediction, compound discovery, and structure–property analysis (Jain et al., 2013; Butler et al., 2018), with advances in reinforcement learning for inverse design (Segler et al., 2018), generative modeling (Granda et al., 2018), and multi-modal integration of simulations and experiments (Pyzer-Knapp et al., 2022). These developments reduce experimental costs and provide interpretable insights, but LLM-based agents remain relatively underexplored in this domain. The inherent complexity of materials science—multiscale interactions, vast compositional spaces, and intricate synthesis pipelines—calls for agent frameworks that unify predictive modeling with experimental feedback. Early systems in autonomous discovery (Gregoire et al., 2023; Maurizi et al., 2024) show promise, but important gaps remain in linking reasoning, planning, and uncertainty quantification. Our work positions materials science as a natural testbed for stress-testing multi-agent reasoning frameworks such as Roundtable Policy, bridging structured evaluation and open-ended scientific narrative.
6 Conclusion & Future Directions
We presented the Roundtable Policy, a inference-time reasoning framework that fuses outputs from multiple LLMs through confidence-weighted aggregation without finetuning base models. Unlike single-model prompting or naive ensemble methods such as majority voting, our approach does not treat all models equally nor rely on opaque emergent convergence, as in debate-style methods. Instead, it introduces a principled mechanism that explicitly encodes historical reliability and uncertainty, producing consensus answers that are not only more accurate but also transparent and auditable. Across structured MultiTask benchmarks and open-ended SingleTask proposal writing, we showed that Roundtable Policy consistently improves factual reliability, reasoning consistency, and narrative coherence, highlighting its value as a trustworthy tool for scientific reasoning.
Looking forward, several avenues can extend this framework. First, adaptive arbitrator strategies could allow the confidence-weight table to evolve dynamically with task complexity, reallocating trust in real time. Second, coupling the fusion process with reinforcement learning and human or experimental feedback may enable autonomous optimization of weight assignments through trial and error. Third, incorporating richer feedback signals—such as calibrated uncertainty, tool-augmented checks, or domain-specific rubrics—could further strengthen robustness in complex scientific workflows.
More broadly, the principles underlying Roundtable Policy generalize beyond the settings studied here. By tailoring consensus mechanisms to domain-specific evaluation criteria, the framework can support applications ranging from drug discovery and computational chemistry to astrophysics and other data-intensive sciences. In this sense, Roundtable Policy is not merely another ensemble method, but a scalable and interpretable paradigm for collective intelligence in LLMs—one that advances the rigor, reliability, and accountability of AI-assisted scientific discovery.
Ethics Statement
Our method significantly enhances the performance of sequential model editing, making it invaluable for updating and managing knowledge in real-world applications. While the ability to directly modify stored knowledge introduces potential risks, such as the introduction of false or harmful information, we strongly urge researchers to implement strict validation and oversight to ensure the ethical use of these techniques. Nevertheless, the original goal of model editing is positive, aiming to facilitate efficient updates of large models in the future. Therefore, we encourage researchers to leverage this technology responsibly and with care.
Reproducibility
To ensure the reproducibility of our findings, detailed implementation instructions for Roundtable Policy can be found in Appendix. Additionally, the source code is publicly available at https://ai4scidisc.github.io/roundtable/. These measures are intended to facilitate the verification and replication of our results by other researchers in the field.
Acknowledement
We gratefully acknowledge the MIT SuperCloud and Texas Advanced Computing Center (TACC) at The University of Texas at Austin for providing computational resources that have contributed to the research results reported within this paper. JL acknowledges support by the Defense Advanced Research Projects Agency (DARPA) under Agreement No. HR00112490369. The content is solely the responsibility of the authors and does not necessarily represent the official views of the funding entities.
Appendix A Notation and Definitions
A.1 Symbols and Abbreviations
: the query space from which tasks are sampled.
: dataset consisting of subtasks with known ground truths (e.g., multiple-choice, fill-in-the-blank, short-answer).
: dataset of structured scientific proposal prompts without ground truth, each consisting of three required paragraphs.
: the number of player agents (pretrained LLMs generating candidate responses). In our experiments .
: the number of grader agents forming the AI Committee (evaluators of player responses). In our experiments .
: the -th player agent, producing a response given an input query.
: the -th grader agent, producing an evaluation score and uncertainty tuple .
: the fusion agent, which synthesizes a final consensus answer from all player responses using the confidence-weight table.
: the query (or query batch) at round .
: the corresponding reference answers (available only in MultiTask).
: the response generated by player in round .
: the pool of all responses in round .
: the score and uncertainty provided by grader for round .
: the aggregated committee evaluation for round .
: the confidence-weight table, a matrix storing trust weights for each player and subtask/rubric dimension.
: the fused consensus response in round .
| Setting | Player ID | Abbreviation | Full Name | Citation |
MultiTask |
001 | GPT-4o | GPT-4o | OpenAI (2024b) |
| 002 | Grok-2 | Grok-2 | xAI (2024) | |
| 003 | Claude-3.5H | Claude-3.5 Haiku | Anthropic (2024) | |
| 004 | GPT-4.5p | GPT-4.5-preview | OpenAI (2024a) | |
| 005 | DS-chat | DeepSeek-Chat | AI (2024) | |
| 006 | Claude-3.7S | Claude-3.7 Sonnet | Anthropic (2025) | |
| 007 | Llama-3.3 | Llama-3.3 | AI (2025) | |
| 008 | Qwen+ | Qwen-plus | Team (2024) | |
| 009 | Gemini-2.0 | Gemini-2.0 | DeepMind (2024) | |
SingleTask |
001 | GPT-4o | GPT-4o | OpenAI (2024b) |
| 002 | GPT-o3 | GPT-o3 | OpenAI (2024d) | |
| 003 | GPT-4.1 | GPT-4.1 | OpenAI (2023a) | |
| 004 | GPT-4.1n | GPT-4.1-nano | OpenAI (2023b) | |
| 005 | GPT-o3-m | GPT-o3-mini | OpenAI (2024e) | |
| 006 | GPT-4o-m | GPT-4o-mini | OpenAI (2024c) | |
| 007 | Grok-3 | Grok-3 | xAI (2025) | |
| 008 | Claude-3.7S | Claude-3.7 Sonnet | Anthropic (2025) | |
| 009 | DS-chat | DeepSeek-Chat | AI (2024) | |
| Setting | Grader ID | Abbreviation | Full Name | Citation |
MultiTask |
001 | GPT-4o | GPT-4o | OpenAI (2024b) |
| 002 | GPT-4.1 | GPT-4.1 | OpenAI (2023a) | |
| 003 | GPT-o3 | GPT-o3 | OpenAI (2024d) | |
| 004 | GPT-o1 | GPT-o1 | OpenAI (2023c) | |
SingleTask |
001 | GPT-4o | GPT-4o | OpenAI (2024b) |
| 002 | GPT-4.1 | GPT-4.1 | OpenAI (2023a) | |
| 003 | GPT-o3 | GPT-o3 | OpenAI (2024d) | |
| 004 | GPT-o1 | GPT-o1 | OpenAI (2023c) |
A.2 Formal Hypothesis Testing Framework for Wilcoxon Signed-Rank Significance Validity
To rigorously assess whether the Roundtable Policy (RT) provides statistically significant improvements over individual LLM baselines, we apply the non-parametric Wilcoxon signed-rank test to paired, round-level performance data. This test evaluates whether the median of the differences between two matched samples significantly exceeds zero, under the null hypothesis of no improvement.
Notation.
Let denote the set of evaluation rounds and be the set of baseline players.
We define:
: score obtained by RT in round ,
: score obtained by player in the same round,
: relative gain of RT over player in round (as percentage).
Hypothesis Test.
We conduct the one-sided Wilcoxon signed-rank test:
Test Procedure.
Given a sample of paired differences , we perform:
1.
Discard entries with .
2.
Rank the absolute values: .
3.
Compute the signed-rank statistic:
(4) -
4.
Using the null distribution of (or a normal approximation), we calculate a -value.
Statistical significance is established if:
Continuity Correction for Normal Approximation.
Under the null hypothesis , the test statistic approximately follows a normal distribution when the sample size is large (typically ). In this case, we compute a -score using the continuity-corrected normal approximation:
where is the number of non-zero values (i.e., rounds with non-zero paired differences), and is the continuity correction term to account for the discreteness of the rank sum distribution.
The resulting -score is then used to compute a one-sided -value based on the standard normal distribution. This approximation is used in our evaluation whenever exceeds the small-sample regime (), ensuring computational efficiency without sacrificing statistical validity.
A.2.1 Application to MultiTask Evaluation
In the MultiTask evaluation, each task represents a distinct skill domain (e.g., math, geoscience, mechanics). Let be the set of benchmark tasks.
For each player and task , and for each round , we compute the percentage gain:
We test each pair independently. To assess holistic advantage, we also aggregate over tasks:
We consider RT statistically superior to player if:
the Wilcoxon test yields on at least 8 of 9 tasks, or
the aggregate test on satisfies .
A.2.2 Application to SingleTask Evaluation
In the SingleTask evaluation, each round produces rubric-based scores over 9 dimensions (covering three traits for each of Background, Methodology, and Impact).
For each player , round , and dimension , we compute:
We conduct Wilcoxon tests:
Per-dimension: for each independently, to assess fine-grained quality improvements.
Aggregate-level: pool all scores:
and test for overall statistical significance.
Interpretation.
The MultiTask analysis evaluates whether RT consistently improves performance across structurally diverse problems and domains. The SingleTask analysis inspects RT’s advantage across nuanced dimensions of academic writing quality. In both cases, aggregate-level tests offer a robust indicator of RT’s overall statistical superiority over individual LLM baselines.
A.3 Formal Framework for Kendall’s Tau Rank Correlation Analysis
To quantify the agreement between different graders’ rankings, we adopt the Kendall’s Tau rank correlation coefficient (Kendall, 1938), a non-parametric statistic that measures the ordinal association between two rankings.
Mathematical Definition.
Let and denote the rankings of items (players) produced by graders and , respectively. For each unordered pair with , define:
A pair is called concordant if and discordant otherwise. Let and be the numbers of concordant and discordant pairs, respectively. Then Kendall’s Tau is defined as
Here indicates perfect agreement, perfect disagreement, and statistical independence of the rankings.
Application to MultiTask.
In the MultiTask setting, each task yields a separate ranking of player models by each grader. For every round and task , we compute for all pairs of graders using their per-round rankings of the nine players. Averaging over all rounds gives
where is the number of valid rounds for task . This produces a correlation matrix per task, summarizing grader consistency across domains such as mathematics, geoscience, and biology.
Application to SingleTask.
In the SingleTask setting, each round produces rubric-based scores across nine dimensions . For each dimension , graders generate rankings of the players, and we compute for all pairs . Averaging across rounds yields
where is the number of valid rounds for dimension . This results in a correlation matrix for each rubric dimension, capturing grader consistency in evaluating creativity, scientific rigor, and logical coherence across sections of the proposal.
Summary.
In both MultiTask and SingleTask, Kendall’s Tau provides a principled measure of inter-grader consistency. The distinction lies in the evaluation unit: tasks for MultiTask versus rubric dimensions for SingleTask.
Appendix B Detailed Demonstration of Roundtable Policy
To provide a concrete illustration of how Roundtable Policy (RT) operates in practice, we present two representative demonstrations: one in the MultiTask regime (Figure 7) and one in the SingleTask regime (Figure 6). These examples showcase how RT transforms fragmented outputs from multiple LLMs into coherent, factually grounded, and narratively rigorous responses.
MultiTask demonstration.
Figure 7 presents a representative MultiTask round, where nine LLM players respond simultaneously to a heterogeneous set of scientific queries spanning mechanics, thermodynamics, semiconductor physics, mathematics, geoscience, and biology. Individual player outputs illustrate the limitations of single-model reasoning: some provide only partial explanations (e.g., noting that paper clips corrode without addressing stress relaxation), others introduce inaccuracies (e.g., oversimplifying nucleation–growth dynamics), and several fail to maintain consistency across domains (e.g., contradictory accounts of polymer glass transition). These fragmented outputs reflect the heterogeneity of LLM competencies—stronger players excel in certain areas, while weaker ones often produce vague or under-specified responses.
By contrast, the fused response generated by RT demonstrates marked improvements in both factual coverage and reasoning depth. For instance, RT not only describes corrosion of paper clips but also contextualizes long-term degradation and mechanical implications; in the nucleation–growth problem, RT integrates complementary insights into thermodynamic driving forces and kinetic limitations, producing a textbook-quality explanation. Similarly, in the polymer glass transition question, RT combines molecular structure, chain length, and plasticizer effects into a unified account that no single player captured in full.
Committee grading further substantiates these gains: RT exhibits higher factual accuracy and reduced uncertainty intervals across tasks. Improvements are particularly pronounced in symbolic and reasoning-intensive domains such as thermodynamics and mathematics, where RT delivers correct solutions while clarifying intermediate reasoning steps. These results highlight how RT balances correctness and clarity across heterogeneous domains.
SingleTask demonstration.
Figure 6 illustrates the SingleTask regime, where each player generates a three-paragraph scientific proposal (Background, Methodology, Impact). Individual drafts vary widely in creativity, methodological depth, and logical flow. Some emphasize novelty but lack technical detail; others offer rigorous methods but fail to contextualize the problem or articulate downstream impact. Such stylistic and structural inconsistencies exemplify the difficulty of producing a coherent scientific narrative with a single LLM.
In contrast, the fused proposal produced by RT exhibits a markedly stronger narrative arc. The Background situates the research problem within its broader scientific context; the Methodology clearly articulates technical contributions with precision; and the Impact section highlights both scientific significance and societal relevance. Grading results confirm consistent improvements across all rubric dimensions—creativity, scientific rigor, and logical coherence—demonstrating that RT effectively mitigates individual stylistic biases while amplifying collective strengths. Even strong baseline players benefit: while their drafts are often polished in one dimension, RT integrates complementary insights to deliver uniformly high-quality writing.
Summary.
Together, these demonstrations show how Roundtable Policy transforms fragmented or uneven model outputs into coherent, high-quality responses. In the MultiTask regime, RT enhances cross-domain reasoning reliability by integrating complementary knowledge across scientific fields. In the SingleTask regime, RT substantially improves narrative structure and explanatory depth in long-form proposal writing. These examples reinforce the central claim that RT enables LLM ensembles to achieve more factual, rigorous, and narratively coherent outputs in scientific applications.
Appendix C Benchmarking Datasets
C.1 Dataset Annotation
We introduce MatSci–Bench, a heterogeneous benchmark designed to evaluate the scientific reasoning capabilities of large language models (LLMs) in materials science and beyond. The benchmark spans six carefully curated sub-domains: (i) Macro Mechanics, (ii) Micro Mechanics, (iii) Optical & Microscopic, (iv) Semiconductor Physics, (v) Thermodynamical Systems, and (vi) Electrical & Electronic Properties. As visualized in Figure 8, these categories are balanced in both question count and token length, enabling systematic cross-domain evaluation of diverse reasoning skills.
This hierarchical structure compels models to integrate reasoning across physical length scales—from atomic-level diffusion and microstructure evolution to bulk elasticity and device performance—while combining symbolic, conceptual, and numerical reasoning. Each sub-domain is paired with task-specific rubrics assessing both (a) factual accuracy and (b) explanatory depth. To further capture reasoning quality, all model outputs must include a chain-of-thought rationale, which is evaluated through two complementary methods: (i) the automatic Causal Reasoning Index (CRI), and (ii) a human-curated explanation rubric (details in Appendix C.4). These fine-grained metrics allow for robust measurement of model performance and principled ablation studies. Table 4 provides representative examples across four broader scientific areas—Biology, Geoscience, Mathematics, and Materials Science—highlighting domain-specific instruction design and diverse task objectives. Unlike prior benchmarks focused narrowly on short factual queries, our dataset emphasizes heterogeneous scientific tasks requiring numerical derivation, mechanistic explanation, and symbolic fluency.
Materials-science subset.
The materials-science questions are drawn from the 2006–2024 UCLA Materials Science and Engineering Ph.D. qualifying examinations. This provenance ensures originality, rigor, and relevance to authentic scientific reasoning. Each problem is annotated with: (i) a numeric ground-truth answer rounded to three significant digits, (ii) an exact LaTeX expression preserving symbolic fidelity, and (iii) a step-by-step expert-verified solution. All annotations are fully compliant with LaTeX and Markdown standards, ensuring reproducibility and integration with downstream evaluation pipelines.
Subdomain Task Instruction Subtask Math Task 007: Math Janet’s ducks lay 16 eggs per day. She consumes 3 eggs for breakfast, uses 4 eggs for baking muffins, and sells the remaining eggs at $2 each. How much does she earn at the farmers’ market? Display strong arithmetic and logical sequencing skills to compute the final earnings. Geoscience Task 008: Geo. Archaeologists record the top layers first, followed by older, deeper layers to study ancient settlements according to the principle of superposition. Demonstrate the ability to interpret stratigraphic evidence and correctly apply the principle of superposition. Biology Task 009: Bio. Does tranexamic acid reduce desmopressin-induced hyperfibrinolysis? Link pharmacological interventions with clinical outcomes based on empirical data. Materials Science Macro Mechanics Task 001: Mech. 1 What is elasticity? What is strain energy? How do we make use of high elastic modulus materials? Explain fundamental mechanical properties and their practical applications in structural design. Micro Mechanics Task 002: Mech. 2 What are interstitial and substitutional diffusion? How do they differ from self-diffusion, and what drives these processes? Differentiate between diffusion mechanisms and explain the role of chemical potential gradients. Electrical & Electronic Task 003: EE What is the difference between homogeneous and heterogeneous nucleation? Demonstrate understanding of nucleation processes and how surfaces or impurities influence phase transitions. Optical & Microscopic Task 004: Opt. Using characterization techniques, how would you distinguish between a mixture of 90% NaCl + 10% LiCl and a mixture of 90% NaCl + 5% LiCl + 5% KCl? Interpret diffraction patterns or elemental analysis data to differentiate material compositions. Thermodynamical Task 005: Thermo. Explain the laws of thermodynamics both mathematically and conceptually. Articulate both the formal mathematical expressions and the underlying concepts of the thermodynamic laws. Semiconductor Task 006: Semi. What is Schrödinger’s equation? Use it to solve for a 1D wavefunction in a simple potential. Demonstrate a solid understanding of quantum mechanics by providing the correct form of Schrödinger’s equation and a step-by-step solution for a 1D potential problem.
Success Criteria: The arbitrator agent confirms that the confidence score of the multi-agents converges.
C.2 Dataset Comparison
| Domain | Q/K | Tokens |
| Geoscience | 1395 | 109,159 |
| Biology | 1000 | 92,866 |
| Math | 1319 | 246,238 |
| Material Science | 539 | 46,150 |
Table 5 benchmarks MatSci–Bench against a range of prominent scientific QA datasets. Whereas prior benchmarks often emphasize scale—providing thousands of short, loosely categorized questions—MatSci–Bench is designed for conceptual depth, compositional reasoning, and graduate-level rigor. For instance, the materials science subset averages tokens per question, more than double that of typical mathematics datasets, reflecting the structural and semantic complexity encountered in authentic scientific workflows.
To construct this benchmark, we developed a hybrid annotation pipeline that balances automated scalability with human quality control. This process begins with an automatic filtering stage, curating candidate prompts from existing resources such as TheoremQA and SciBench. GPT-4 is then employed to generate positive tool functions—programmatic calls capable of faithfully executing or reasoning over the problem content—which are subsequently audited, refined, or discarded by expert annotators.
To further prevent models from exploiting shallow heuristics or brittle shortcuts, annotators also construct complementary negative tool functions that intentionally fail or only partially match the task. Together, these positive–negative mappings define the final tool interface for each problem, ensuring that evaluation targets robust reasoning rather than superficial pattern recognition. Full implementation details, including annotation consistency and rubric alignment statistics, are reported in Appendix C.4.
C.3 QK-Pair Overview
Unlike ensemble approaches that rely on static averaging or manually crafted heuristics, our framework enables fine-grained, task-dependent variability in response aggregation. It supports dynamic arbitration by adapting both model selection and integration strategies to the empirical strengths of each LLM on a per-task basis, thereby facilitating compositional reasoning across heterogeneous scientific inputs.
Tasks are categorized into two principal types according to the nature of supervision:
Ground-truth-based tasks. This category includes multiple-choice questions, numerical derivations, and factual reasoning problems, where each model’s prediction can be directly compared against a known reference. Scoring in this regime is deterministic and fully automated.
Non-ground-truth tasks. This category includes open-ended generation problems, such as scientific proposal writing, where no unique ground truth exists. Evaluation is therefore conducted by the AI Committee using rubric-based qualitative assessment, accounting for creativity, scientific rigor, and logical coherence.
To ensure reproducibility and modularity, all tasks are stored in a structured format, where denotes the input query and the expected response (if applicable). In ground-truth settings, provides the exact solution or numeric tolerance reference. In non-ground-truth settings, instead encodes evaluation rubrics or stylistic constraints, guiding qualitative assessment rather than exact matching.
We evaluate the framework under two complementary protocols:
MultiTask. Each round presents a batch of heterogeneous tasks spanning multiple scientific domains as shown in Figure 9. The AI Committee jointly evaluates player responses, testing the arbitration system’s ability to generalize across diverse input types, modalities, and difficulty levels.
SingleTask. Each round focuses on a scientific proposal generation as shown in Figure 10, allowing the Committee to learn nuanced arbitration strategies within a fixed domain and structured output format.
Together, these protocols provide a comprehensive evaluation of arbitration dynamics under both breadth-oriented (MultiTask) and depth-oriented (SingleTask) conditions. Importantly, all evaluations are conducted without modifying the internal parameters of the player models, thereby preserving their pretrained competencies while enabling compositional reasoning through learned arbitration.
C.4 MatSci–Bench
Table 6 illustrates the mapping between several UCLA courses (from the prelim exams spanning 2006–2024) and our six defined subdomains in Material Science. The table indicates the number of questions per course and marks (✓) the corresponding subdomains covered by each course. And Table 7 Mapping from numeric course codes to official course names in the UCLA MSE Graduate Program.
In our multi-task framework, each question is paired with a corresponding key—forming a question-key pair—that encapsulates its unique characteristics and domain-specific attributes. This design enables a fine-grained classification of tasks, as each key serves as an identifier for the subdomain and the underlying problem-solving facet it represents. In our dataset, which is constructed from high-quality, authoritative UCLA prelim exams, each question is not only rigorously structured but also assigned a key that maps to distinct subdomains within Material Science. This approach allows for a combinatorial evaluation across various tasks and provides a nuanced understanding of the performance of large language models (LLMs). Moreover, the explicit association between questions and keys facilitates a higher resolution in assessment, offering insights into the model’s capability to address the subtle distinctions inherent in each category.
| Course # | Mechanism | Property | ||||
| Macro | Micro |
Electrical &
Electronic |
Optical &
Microscopic |
Semicon-
ductor |
Thermo-
dynamical |
| 200 | ✓ | ✓ | ||||
| 201 | ✓ | ✓ | ||||
| 202 | ✓ | ✓ | ||||
| 210 | ✓ | ✓ | ||||
| C211 | ✓ | ✓ | ||||
| C212 | ✓ | ✓ | ||||
| M213 | ✓ | ✓ | ||||
| 213L | ✓ | ✓ | ||||
| 214 | ✓ | ✓ | ||||
| 216 | ✓ | ✓ | ||||
| 221 | ✓ | ✓ | ✓ | |||
| 222 | ✓ | ✓ | ||||
| 223 | ✓ | ✓ | ||||
| 224 | ✓ | ✓ | ||||
| 225 | ✓ | ✓ | ||||
| 226 | ✓ | ✓ | ||||
| 243A | ✓ | ✓ | ||||
| 243C | ✓ | ✓ | ✓ | |||
| 246A | ✓ | ✓ | ||||
| 246B | ✓ | ✓ | ||||
| 246D | ✓ | ✓ | ||||
| 247 | ✓ | ✓ | ||||
| 248 | ✓ | ✓ | ||||
| 250B | ✓ | ✓ | ||||
| 251 | ✓ | ✓ | ||||
| 252 | ✓ | ✓ | ||||
| 253 | ✓ | ✓ | ||||
| 261 | ✓ | ✓ | ||||
| 262 | ✓ | ✓ | ||||
| CM263 | ✓ | ✓ | ||||
| 270 | ✓ | ✓ | ||||
| 271 | ✓ | ✓ | ||||
| 272 | ✓ | ✓ | ||||
| CM280 | ✓ | ✓ | ||||
| 282 | ✓ | ✓ | ||||
| 296 | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
| M297B | ✓ | ✓ | ||||
| M297C | ✓ | ✓ | ||||
| 298 | ✓ | ✓ | ||||
| Course # | Course Name |
| 200 | Principles of Materials Science I |
| 201 | Principles of Materials Science II |
| 202 | Thermodynamics of Materials |
| 210 | Diffraction Methods in Science of Materials |
| C211 | Introduction to Materials Characterization B (Electron Microscopy) |
| C212 | Cultural Materials Science II: Characterization Methods in Conservation of Materials |
| M213 | Cultural Materials Science I: Analytical Imaging and Documentation in Conservation of Materials |
| 213L | Cultural Materials Science Laboratory: Technical Study |
| 214 | Structure, Properties, and Deterioration of Materials: Rock Art, Wall Paintings, Mosaics |
| 216 | Science of Conservation Materials and Methods I |
| 221 | Science of Electronic Materials |
| 222 | Growth and Processing of Electronic Materials |
| 223 | Materials Science of Thin Films |
| 224 | Deposition Technologies and Their Applications |
| 225 | Materials Science of Surfaces |
| 226 | Si-CMOS Technology: Selected Topics in Materials Science |
| 243A | Fracture of Structural Materials |
| 243C | Dislocations and Strengthening Mechanisms in Solids |
| 246A | Mechanical Properties of Nonmetallic Crystalline Solids |
| 246B | Structure and Properties of Glass |
| 246D | Electronic and Optical Properties of Ceramics |
| 247 | Nanoscale Materials: Challenges and Opportunities |
| 248 | Materials and Physics of Solar Cells |
| 250B | Advanced Composite Materials |
| 251 | Chemistry of Soft Materials |
| 252 | Organic Polymer Electronic Materials |
| 253 | Bioinspired Materials |
| 261 | Risk Analysis for Engineers and Scientists |
| 262 | Probabilistic Modeling and Simulation of Complex Systems |
| CM263 | Electrochemical Processes |
| 270 | Computer Simulations of Materials |
| 271 | Electronic Structure of Materials |
| 272 | Theory of Nanomaterials |
| CM280 | Introduction to Biomaterials |
| 282 | Exploration of Advanced Topics in Materials Science and Engineering |
| 296 | Seminar: Advanced Topics in Materials Science and Engineering |
| M297B | Material Processing in Manufacturing |
| M297C | Composites Manufacturing |
| 298 | Seminar: Engineering |
Appendix D Fine-grained Results
D.1 MultiTask
Player ID Task ID # (%) (%) p-value Significance Overall Player_001 GPT-4o 001 36000 3.664534 29.115645 ✓ ✓ 002 36000 9.990613 66.756230 ✓ 003 36000 4.418581 32.628933 ✓ 004 36000 18.290543 103.900081 ✓ 005 36000 3.507243 49.094848 ✓ 006 36000 4.547669 29.876047 ✓ 007 36000 18.619909 110.561195 ✓ 008 36000 4.205786 76.940468 ✓ 009 36000 -13.817889 121.279772 ✗ Player_002 Grok-3 001 36000 2.442676 26.869884 ✓ ✓ 002 36000 11.658807 87.033589 ✓ 003 36000 2.922847 25.540567 ✓ 004 36000 10.040225 97.107301 ✓ 005 36000 2.754232 68.334607 ✓ 006 36000 7.795831 50.756077 ✓ 007 36000 7.037656 73.901148 ✓ 008 36000 12.159710 90.508914 ✓ 009 36000 -14.196161 124.567204 ✗ Player_003 Claude-3.5Haiku 001 36000 9.356006 49.469078 ✓ ✓ 002 36000 25.634084 100.488313 ✓ 003 36000 8.903574 52.492393 ✓ 004 36000 24.974510 120.272436 ✓ 005 36000 18.091843 89.503620 ✓ 006 36000 19.206568 71.695176 ✓ 007 36000 29.644628 127.261225 ✓ 008 36000 22.433517 119.787354 ✓ 009 36000 -8.057303 138.391228 ✗ Player_004 GPT-4.5-preview 001 36000 26.864778 50.638001 ✓ ✓ 002 36000 24.364182 76.170729 ✓ 003 36000 28.888756 93.939402 ✓ 004 36000 18.305048 104.942899 ✓ 005 36000 18.611880 82.101495 ✓ 006 36000 17.971459 45.269719 ✓ 007 36000 32.007776 136.936369 ✓ 008 36000 5.146633 83.614309 ✓ 009 36000 -11.464038 128.061246 ✗ Player_005 Deepseek-chat 001 36000 15.448390 50.327417 ✓ ✓ 002 36000 21.685092 88.100774 ✓ 003 36000 11.772803 42.894447 ✓ 004 36000 14.794087 99.773548 ✓ 005 36000 9.123905 59.031496 ✓ 006 36000 12.303097 45.278523 ✓ 007 36000 52.148935 167.006012 ✓ 008 36000 11.541473 99.093907 ✓ 009 36000 -8.504933 137.294744 ✗
Player ID Task ID # (%) (%) p-value Significance Overall Player_006 Claude-3.7Sonnet 001 36000 10.483677 100.447956 ✓ ✓ 002 36000 20.034584 140.497283 ✓ 003 36000 7.986260 57.805017 ✓ 004 36000 11.881086 139.264504 ✗ 005 36000 4.431706 106.492570 ✓ 006 36000 11.670759 84.754451 ✓ 007 36000 65.488159 182.402404 ✓ 008 36000 10.358145 91.889615 ✓ 009 36000 -16.646343 131.246587 ✗ Player_007 Llama-3.3 001 36000 26.016269 81.867305 ✓ ✓ 002 36000 43.795534 133.423412 ✓ 003 36000 29.847791 118.361082 ✓ 004 36000 33.835778 146.064900 ✓ 005 36000 31.906208 125.420480 ✓ 006 36000 29.654251 95.808593 ✓ 007 36000 12.560069 113.533051 ✓ 008 36000 24.151429 138.741394 ✓ 009 36000 -8.214252 139.363305 ✗ Player_008 Qwen-plus 001 36000 12.702783 53.480581 ✓ ✓ 002 36000 24.671245 105.447534 ✓ 003 36000 12.217150 54.159535 ✓ 004 36000 19.478679 117.267431 ✓ 005 36000 6.452739 58.406561 ✓ 006 36000 18.500964 66.688703 ✓ 007 36000 16.654888 101.522057 ✓ 008 36000 15.475509 98.223206 ✓ 009 36000 -13.418640 129.270826 ✗ Player_009 Gemini-2.0 001 36000 48.597322 101.645241 ✓ ✓ 002 36000 56.207001 136.978257 ✓ 003 36000 30.807297 86.405085 ✓ 004 36000 40.730016 138.029856 ✓ 005 36000 39.623737 120.151633 ✓ 006 36000 43.606884 96.069039 ✓ 007 36000 127.050040 253.842133 ✓ 008 36000 19.718319 112.418143 ✓ 009 36000 15.955998 173.200269 ✓
D.2 SingleTask
Player ID Paragraph Dimension # (%) (%) p-value Significance Overall Player_001 GPT-4o Background Creativity 36000 37.150702 49.336875 ✓ ✓ Scientific Rigor 36000 30.257828 46.850598 ✓ Logical Coherence 36000 11.733576 16.004795 ✓ Methodology Creativity 36000 24.931209 31.921273 ✓ Scientific Rigor 36000 31.838154 45.990879 ✓ Logical Coherence 36000 17.282977 21.020188 ✓ Impact Creativity 36000 32.606157 40.063019 ✓ Scientific Rigor 36000 41.271411 57.887977 ✓ Logical Coherence 36000 16.718015 19.804013 ✓ Player_002 GPT-o3 Background Creativity 36000 12.164463 112.053304 ✓ ✓ Scientific Rigor 36000 13.556005 141.849680 ✓ Logical Coherence 36000 7.902093 88.022126 ✓ Methodology Creativity 36000 12.242360 162.582687 ✓ Scientific Rigor 36000 13.742029 176.441027 ✓ Logical Coherence 36000 12.246663 149.530625 ✓ Impact Creativity 36000 14.822979 162.636940 ✓ Scientific Rigor 36000 19.794367 172.889406 ✓ Logical Coherence 36000 13.375266 149.888340 ✓ Player_003 GPT-4.1 Background Creativity 36000 23.490292 76.488563 ✓ ✓ Scientific Rigor 36000 15.516892 73.369125 ✓ Logical Coherence 36000 7.636895 70.613455 ✓ Methodology Creativity 36000 13.841302 73.356046 ✓ Scientific Rigor 36000 13.786087 74.585712 ✓ Logical Coherence 36000 9.985966 70.810643 ✓ Impact Creativity 36000 18.657843 73.142286 ✓ Scientific Rigor 36000 22.049522 75.719467 ✓ Logical Coherence 36000 11.006198 72.520043 ✓ Player_004 GPT-4.1-nano Background Creativity 36000 22.875212 41.517104 ✓ ✓ Scientific Rigor 36000 16.303948 31.393619 ✓ Logical Coherence 36000 7.715497 14.464709 ✓ Methodology Creativity 36000 14.951288 21.488109 ✓ Scientific Rigor 36000 14.777490 25.696003 ✓ Logical Coherence 36000 10.347582 16.738338 ✓ Impact Creativity 36000 19.842341 27.921968 ✓ Scientific Rigor 36000 23.655715 38.425349 ✓ Logical Coherence 36000 11.784470 18.251267 ✓ Player_005 GPT-o3-mini Background Creativity 36000 12.773348 27.212152 ✓ ✓ Scientific Rigor 36000 9.060914 24.819729 ✓ Logical Coherence 36000 5.406290 14.736602 ✓ Methodology Creativity 36000 10.166683 16.487638 ✓ Scientific Rigor 36000 7.171348 19.411696 ✓ Logical Coherence 36000 6.591953 13.835775 ✓ Impact Creativity 36000 13.300165 22.726699 ✓ Scientific Rigor 36000 14.308252 31.344613 ✓ Logical Coherence 36000 8.408882 16.386089 ✓
Player ID Paragraph Dimension # (%) (%) p-value Significance Overall Player_006 GPT-4O-mini Background Creativity 36000 33.709469 47.064151 ✓ ✓ Scientific Rigor 36000 32.286309 53.299864 ✓ Logical Coherence 36000 13.247615 20.226189 ✓ Methodology Creativity 36000 25.170102 32.705104 ✓ Scientific Rigor 36000 28.879875 45.066162 ✓ Logical Coherence 36000 17.627694 24.373974 ✓ Impact Creativity 36000 32.096396 42.641222 ✓ Scientific Rigor 36000 42.631153 63.716542 ✓ Logical Coherence 36000 19.211624 26.269863 ✓ Player_007 Grok-3 Background Creativity 36000 21.453687 36.436965 ✓ ✓ Scientific Rigor 36000 20.427351 37.117935 ✓ Logical Coherence 36000 8.073442 15.451901 ✓ Methodology Creativity 36000 14.424544 21.979675 ✓ Scientific Rigor 36000 16.424389 26.740515 ✓ Logical Coherence 36000 10.934741 16.722313 ✓ Impact Creativity 36000 18.844803 28.521192 ✓ Scientific Rigor 36000 25.975916 41.628876 ✓ Logical Coherence 36000 11.808618 18.702339 ✓ Player_008 Claude-3.7Sonnet Background Creativity 36000 12.640859 25.792853 ✓ ✓ Scientific Rigor 36000 7.857101 17.472430 ✓ Logical Coherence 36000 3.793134 9.710774 ✓ Methodology Creativity 36000 6.475117 11.493763 ✓ Scientific Rigor 36000 6.547614 14.204805 ✓ Logical Coherence 36000 5.655090 10.774578 ✓ Impact Creativity 36000 8.043748 15.478956 ✓ Scientific Rigor 36000 12.470529 24.399598 ✓ Logical Coherence 36000 6.367473 13.337112 ✓ Player_009 Deepseek-Chat Background Creativity 36000 7.791205 8.969878 ✓ ✓ Scientific Rigor 36000 7.757954 8.056653 ✓ Logical Coherence 36000 3.207566 5.639610 ✓ Methodology Creativity 36000 7.612794 7.235925 ✓ Scientific Rigor 36000 8.278456 7.740058 ✓ Logical Coherence 36000 6.056968 5.893318 ✓ Impact Creativity 36000 8.014463 7.344285 ✓ Scientific Rigor 36000 10.484719 7.746186 ✓ Logical Coherence 36000 5.287961 5.505316 ✓
D.3 AI Graders
A central component of our evaluation pipeline is the use of AI graders (arbitrators) to provide rankings of model outputs. Unlike human experts, who bring domain expertise but are costly and inconsistent across annotators, AI graders offer scalable, replicable, and transparent evaluation. Nevertheless, just as in human peer review, the presence of bias (systematic preference of a grader) and the degree of consistency (agreement across graders) are critical for ensuring the reliability of aggregated judgments. The purpose of this section is to examine these properties in detail, thereby motivating our design choice of adopting multiple independent graders rather than relying on a single arbitrator.
Motivation.
Bias in grading arises when one arbitrator disproportionately favors certain ranks across repeated evaluations. For example, a grader may consistently assign lower (more critical) ranks, while another may tend to be more lenient. If such biases are not controlled, downstream evaluations may unfairly penalize or reward certain systems, obscuring the true effect of the Roundtable Policy. Consistency, on the other hand, reflects whether different graders produce comparable relative orderings. High consistency increases our confidence that the rankings reflect robust quality differences, while low consistency signals noise and arbitrariness. By quantifying both bias and consistency, we establish the validity of our evaluation methodology.
Methodology.
We assess grader bias by visualizing the empirical rank distributions assigned by each arbitrator. Figure 3 presents kernel density estimates of the rank distributions for four graders across both the SingleTask and MultiTask evaluations. Systematic deviations in the distributions (e.g., skewed toward higher or lower ranks) indicate bias. To measure consistency, we compute pairwise Kendall’s Tau correlations between graders’ rankings. This non-parametric statistic captures the degree of ordinal agreement, with values close to indicating strong concordance, values near suggesting randomness, and negative values reflecting disagreement. The results are shown as inset heatmaps in Figure 3.
Findings.
In the SingleTask setting, we observe substantial variation in grader behavior. The rank distributions are uneven, with some graders systematically harsher or more generous than others. The Kendall’s Tau correlations confirm this observation, with moderate agreement values that indicate significant variability across graders. Such findings echo common phenomena in human peer review, where individual reviewers bring idiosyncratic standards and biases. By contrast, in the MultiTask setting, the distributions are more balanced, and the Kendall’s Tau values are consistently higher. This suggests that when grading across diverse tasks, individual biases are averaged out, leading to more stable and reliable inter-grader consensus.
Implications.
These analyses reinforce two important principles. First, reliance on a single grader is inherently risky due to potential bias; employing multiple graders and aggregating their judgments is essential for fairness. Second, broader task coverage increases reliability, as MultiTask grading mitigates the influence of any single arbitrator’s bias. Together, these results highlight the necessity of designing evaluation protocols that balance scalability with robustness, ensuring that AI grader-based assessments approximate the fairness and credibility traditionally associated with human expert panels.

### v2

@@
-Roundtable Policy: Improving Scientific Reasoning and Narratives through Confidence-Weighted Consensus of LLMs
+Roundtable Policy: Confidence-Weighted-Consensus Aggregation Improves Multi-Agent-System Reasoning
 Abstract
-Large language models (LLMs) have demonstrated remarkable capabilities not only in language generation but also in advancing scientific discovery.
-A growing body of work has explored ways to improve their reasoning, from self-consistency and chain-of-thought to multi-agent debate.
-Inspired by the dynamics of scientific committees and The Society of Mind, we introduce Roundtable Policy, a complementary inference-time reasoning framework that performs inference through the weighted consensus of multiple LLMs.
-Our findings indicate that this approach significantly enhances reasoning in complex heterogeneous scientific tasks and improves scientific narratives in terms of creativity, rigor, and logical coherence, while reducing hallucinations that single models are prone to.
-Our approach emphasizes structured and interpretable consensus rather than opaque convergence, while requires only black-box access and uniform procedures, making it broadly applicable to multi-LLM reasoning.
-Project website at https://ai4scidisc.github.io/roundtable/.
+Multi-agent systems have demonstrated exceptional performance in downstream tasks beyond diverse single agent baselines.
+A growing body of work has explored ways to improve their reasoning and collaboration, from vote, debate, to complex interaction protocols.
+However, it still remains opaque why specific choice would be preferred in multi-agent systems.
+Inspired by the decision-making mechanism of democratic committees and The Society of Mind, we introduce Roundtable Policy, an inference-time reasoning framework for multi-agent systems that performs inference through the weighted consensus of multiple LLMs.
+Through extensive experiments, we demonstrate its that this approach significantly enhances reasoning in complex heterogeneous scientific tasks.
+Roundtable Policy emphasizes structured and interpretable inference rather than opaque convergence, while requires only black-box access and uniform procedures, making it broadly applicable to diverse multi-agent systems.
 1 Introduction
-Large language models (LLMs) have demonstrated impressive capabilities in text generation, reasoning, and few-shot adaptation.
-These advances have opened exciting opportunities for scientific discovery, where complex, interdisciplinary problems demand not only accurate reasoning but also coherent narrative construction.
-However, current LLMs may hallucinate facts, introduce implausible logical jumps, or produce fragmented explanations.
-Such limitations are particularly problematic in scientific contexts, where reliability and clarity are equally critical.
-For instance, a scientific proposal must not only provide factually correct content, but also organize it into a structured and persuasive narrative.
-Addressing these dual requirements calls for mechanisms that go beyond the capacity of a single LLM.
-A natural motivation is therefore to leverage multiple LLMs, aggregating their complementary strengths.
-Prior work has explored self-consistency with multiple sampled reasoning paths (Wang et al., 2022), multi-agent debate frameworks (Du & et al., 2023; Madaan & et al., 2023), and tool-augmented multi-agent systems (Schick et al., 2023; Ghafarollahi & Buehler, 2024).
-However, these approaches are often limited to single-task reasoning, rely on emergent dynamics that are difficult to interpret, or lack systematic mechanisms for weighting contributions across heterogeneous domains.
-Ensemble-style methods such as majority voting treat all models equally, overlooking that different models excel in different domains or evaluation dimensions.
-Debate-style methods, while sometimes converging to better outputs, remain inherently opaque: it is often unclear why a particular answer was selected, or hard to quantify which models were more trustworthy in the process.
-In scientific applications, this lack of transparency poses a major barrier, since researchers require not only correct outputs but also interpretable evidence supporting the decision.
-At a broader level, scientific problem solving is inherently multi-faceted, requiring both precise reasoning and coherent long-form narratives.
-While individual LLMs have made impressive progress, they still face challenges in consistently meeting these diverse demands across domains and tasks.
-This motivates the need for a structured approach to harness collective intelligence—one that integrates complementary expertise, surfaces and resolves disagreements transparently, and preserves accountability in the decision process.
-Building on these considerations, we introduce Roundtable Policy, an inference-time framework that formalizes deliberation among multiple LLMs into a principled process for trustworthy scientific reasoning and narrative generation (Figure 1).
-In our approach, diverse models first propose candidate responses, which are then jointly evaluated to record their performances.
-These evaluations are distilled into a confidence-weight table that encodes both historical reliability and uncertainty for each candidate across tasks and dimensions.
-At inference, this confidence-weight table guides a fusion process that integrates the diverse responses into a single consensus answer—balancing factual accuracy with narrative coherence while preserving transparency in how the decision is reached.
-As illustrated in Figure 1, even on a scientific question such as explaining why nucleation accelerates at greater undercooling while growth dominates at smaller undercooling, individual models provide partial or inconsistent accounts.
-The roundtable process fuses these perspectives into a coherent explanation, showing how consensus-driven reasoning can elevate fragmented insights into a more complete scientific narrative.
-We evaluate Roundtable Policy in two complementary scientific regimes.
-In the MultiTask setting, heterogeneous problems from geoscience, biology, mathematics, and materials science are solved simultaneously, testing cross-domain reasoning.
-In the SingleTask setting, models generate structured scientific proposals, testing long-form narrative construction.
-Across both, our findings show that Roundtable Policy consistently improves factual reliability, reasoning quality, and narrative coherence, yielding an average 13.01% improvement in MultiTask and 11.04% improvement in SingleTask compared to individual models.
-Unlike ensemble-style methods such as majority voting, which treat all models equally, our approach adaptively weights models based on demonstrated reliability.
-And in contrast to debate-style frameworks, whose convergence dynamics are opaque, Roundtable Policy produces explicit, auditable signals that make the aggregation process transparent.
-This combination of principled weighting and interpretability distinguishes our method as a scalable and accountable alternative for multi-model scientific reasoning.
-In summary, our contributions are threefold: (i) we propose Roundtable Policy, an inference-time framework for performance-weighted consensus aggregation among multiple LLMs; (ii) we demonstrate that this approach enhances both factual reasoning and scientific narrative across heterogeneous tasks; and (iii) we release a graduate-level materials science benchmark derived from UCLA PhD qualifying exams, which serves as a rigorous, domain-authentic testbed and complements existing evaluations of scientific LLMs.
-More broadly, the principles underlying Roundtable Policy apply across multi-model scientific workflows, offering an interpretable and generalizable path toward AI-assisted reasoning and communication in science.
+Scientific reasoning often requires synthesizing multiple, partially reliable perspectives.
+This process needs not only predicting a single correct answer, but also forming a coherent and reliable consensus across partially correct, incomplete, or even conflicting explanations.
+Large language models (LLMs) have shown remarkable capabilities in generating informative responses, motivating their use in scientific discovery and reasoning tasks.
+However, when applied to answering complex scientific problems, individual models may hallucinate facts, make brittle logical jumps, or emphasize different aspects of a problem.
+These behaviors reflect not only limitations of individual models, but also a broader challenge: scientific reasoning is not fully captured by isolated, single-shot predictions, but instead requires mechanisms for aggregating multiple uncertain and heterogeneous reasoning paths.
+For instance, to write a scientific proposal must not only provide factually correct content, but also organize it into a structured and persuasive narrative.
+Addressing these requirements calls for mechanisms that can reconcile multiple reasoning paths.
+A natural motivation is therefore to leverage multi-agent systems, collecting their complementary strengths via diverse reasoning trajectories.
+Prior work has explored from majority (Wang et al., 2023), debate (Du et al., 2024; Liang et al., 2024), to other ensemble approaches (Chen et al., 2024; Wang et al., 2025).
+While effective in many cases, these approaches exhibit inherent limitations: majority vote, treat all models equally, overlooking that different models excel in different domains; Debate-style methods, while sometimes converging to better outputs, remain inherently opaque: it is unclear why a particular answer was selected, or hard to quantify which models were more trustworthy in the process.
+These observations all point to a central question: how can we effectively combine complementary strengths across models to solve complex scientific tasks?
+From this perspective, multi-agent reasoning can be viewed as a problem of consensus formation across heterogeneous reasoning trajectories.
+We instantiate this principle with Roundtable Policy, an inference-time reasoning framework that explicitly models agent reliability with uncertainty across task-specific dimensions (Figure 2), producing auditable signals to make inference more transparent and interpretable without modifying base models.
+In our framework, diverse agents first propose candidate responses, which are evaluated and recorded across tasks.
+Then these historical evaluations are distilled into a confidence-weight table that captures both agent reliability and uncertainty.
+During inference, this structured memory guides aggregation, enabling the system to integrate diverse perspectives into a coherent consensus while remaining transparent about how trust is allocated across agents.
+This perspective also motivates how scientific reasoning systems should be evaluated.
+Current evaluation has largely relied on benchmarks with isolated, singe-domain tasks.
+Representative examples such as GSM8K (Cobbe et al., 2021), PubMedQA (Jin et al., 2019), and AlphaGeometry (Trinh et al., 2024) have been instrumental in assessing domain-specific knowledge and pointwise accuracy.
+However, these benchmarks typically assume short-context, single-problem settings, and therefore fail to capture key characteristics of scientific contexts, which often require integrating knowledge across domains and maintaining logical consistency over long contexts.
+For example, writing a scientific proposal involves synthesizing background knowledge, justifying methodologies, and keeping long-context consistency, rather than solving an isolated arithmetic problem like 4+23*6+24-24*12 (GSM8K).
+This mismatch becomes sharper in multi-agent settings, where reasoning quality also depends on how complementary strengths are aggregated.
+To address this gap, we design two complementary benchmarks that capture distinct but essential aspects of scientific reasoning: ScienceEval and ScienceNarrative.
+ScienceEval evaluates cross-domain reasoning by jointly querying agents on heterogeneous problems spanning geoscience (Deng et al., 2024), biology (Jin et al., 2019), mathematics (Cobbe et al., 2021), and multi-faceted physical science (§C).
+ScienceNarrative focuses on long-context reasoning consistency by requiring agents to generate structured scientific proposals under contextual templates.
+Across both benchmarks, Roundtable Policy consistently improves factual reliability and narrative coherence, achieving average gains of 13.01% in ScienceEval and 11.04% in ScienceNarrative over diverse baseline agents.
+In summary, our contributions are threefold.
+(1) We formulate multi-agent reasoning as a problem of consensus formation across heterogeneous reasoning trajectories, and instantiate this perspective with Roundtable Policy, an inference-time reasoning framework.
+(2) We introduce two complementary benchmarks, ScienceEval and ScienceNarrative, to evaluate cross-domain and long-context consistency and accuracy, providing domain-authentic testbeds aligned with the demands of complex scientific reasoning.
+(3) We demonstrate the efficacy of Roundtable Policy through extensive experiments.
 2 Methodology
-This section formalizes the Roundtable Policy framework.
-We begin with the intuition behind the Roundtable Policy and then present the task settings and notation (a complete list of symbols is provided in Appendix §A).
-We next describe how the confidence-weight table is learned from repeated interactions between players and graders, and how it subsequently guides inference through consensus-weighted fusion.
+In this section, we begin with the intuition behind the Roundtable Policy and then describe the operational mechanism of this framework.
 2.1 Design Intuitions behind Roundtable Policy
-Scientific decision-making rarely relies on a single voice.
-In research committees, for example, each member independently evaluates a proposal, but the final judgment is not determined by a simple vote.
-Instead, reviewers’ assessments are weighed by domain expertise, reliability, and the clarity of justification before being synthesized into a collective decision.
-This process balances agreement with accountability, ensuring that the outcome is both credible and transparent.
-Roundtable Policy draws on the same principle.
-Rather than treating LLMs as isolated predictors, we view them as members of a committee.
-A set of players propose candidate responses informed by diverse inductive biases.
-These are evaluated by an ensemble of graders, whose rubric-based assessments provide explicit signals of quality and uncertainty.
-Finally, a structured fusion step aggregates these evaluations into a single consensus output—analogous to how a committee chair synthesizes diverse reviews into a reasoned decision.
-This design is further motivated by the perspective of The Society of Mind (Minsky, 1988), which frames intelligence as the coordination of many specialized, relatively simple agents.
-By embracing this philosophy, Roundtable Policy transforms the inherent diversity of LLMs into a strength, enabling collective reasoning that is both interpretable and robust.
-2.2 Task Settings
-We evaluate Roundtable Policy in two complementary scientific regimes that mirror the dual demands of scientific reasoning: solving pointwise problems and constructing coherent narratives.
-MultiTask.
-Models are simultaneously queried with nine heterogeneous scientific questions spanning geoscience, biology, mathematics, and materials science (with six distinct sub-categories).
-Task formats include multiple-choice, fill-in-the-blank, and short-answer questions.
-Intuitively, this setting resembles a written exam with diverse problem types—success requires breadth of expertise rather than mastery of a single domain.
-SingleTask.
-Models are prompted to generate structured scientific proposals following a standardized three-paragraph template with sections for Background, Methodology, and Impact.
-This setting captures the narrative aspect of scientific work: beyond producing isolated facts, one must weave them into a logically consistent and persuasive storyline.
-Across both regimes, we deploy pretrained LLMs (players) to generate candidate responses and LLMs (graders) to form the AI Committee.
-Just as in a real committee, players provide diverse answers while graders act as evaluators who judge each contribution against shared standards.
-In MultiTask, graders jointly evaluate the responses per round, analogous to examiners scoring a large pool of student answers.
-In SingleTask, graders assess proposals paragraph-by-paragraph along three rubric dimensions—creativity, scientific rigor, and logical coherence—reflecting the multifaceted criteria by which scientific proposals are judged.
-Each evaluation yields a trust weight with two components: (i) a quality score in , where negative values indicate incorrect or misleading responses, and (ii) a calibrated 95% confidence interval representing grader certainty.
+In multi-agent reasoning, agents often differ in expertise and reliability across tasks.
+Effective collaboration therefore requires inference-time mechanisms that can account for these differences, rather than treating all agents or reasoning paths as equally trustworthy.
+A useful intuition comes from democratic committees.
+Members independently evaluate a proposal, but final decisions are rarely made by simple voting.
+Instead, assessments are weighted by expertise, past reliability, and confidence, leading to decisions that are both robust and interpretable.
+Roundtable Policy follows this intuition by viewing multi-agent reasoning as a committee-style inference process: agents propose diverse reasoning trajectories, their task-specific reliability is assessed over time, and a final consensus is formed by weighting historical reliability and confidence of each member.
+2.2 Multi-Phase Setting
+To realize this intuition, we structure multi-agent reasoning into multiple phases that explicitly separate proposal, evaluation, and consensus formation. Notations follow §A.
+Proposition Phase.
+In the proposition phase, a set of player (or proposer) agents() independently generate candidate responses to the same query.
+This design encourages diversity in reasoning paths and solution strategies, providing a rich pool of alternative perspectives for subsequent aggregation.
+Inference Phase.
+In the inference phase, an aggregator (or fusion) agent() synthesizes the candidate responses produced by the players.
+The aggregation is explicitly conditioned on each player’s historical performance profile, represented as a score with associated uncertainty.
+By incorporating this structured reliability information, the aggregator can weigh contributions beyond instantaneous agreement, enabling informed consensus formation.
+Backward Phase.
+In the backward phase, a committee of LLM-based grader agents() evaluates the players’ responses under task-specific or rubric-based criteria.
+These evaluations yield quantitative assessments of both quality and uncertainty, which are used to update the historical performance records of each player.
+Importantly, this phase updates only the external confidence-weight table and does not modify the parameters of any underlying language model.
+2.3 Encoding Historical Performance via Confidence-Weight Table
+In the backward phase, each evaluation yields a trust weight with two components: (i) a quality score in , where negative values indicate incorrect or misleading responses, and (ii) a calibrated 95% confidence interval representing the uncertainty.
 Together, these form the equivalent of both a “grade” and a measure of “grading confidence,” ensuring that downstream decisions reflect not only evaluation outcomes but also their reliability.
-Notation follows Appendix §A denotes the -th player agent, the -th grader, and the fusion agent.
-Player responses and grader scores are indexed by round .
-2.3 Learning the Confidence-Weight Table
-The confidence-weight table serves as a structured memory of each player’s historical reliability across subtasks (MultiTask) or rubric dimensions (SingleTask).
-Unlike fine-tuning approaches, Roundtable Policy does not alter the parameters of the underlying LLMs.
-Instead, is iteratively updated from repeated rounds of evaluation, much like how a committee gradually learns which members are most reliable in which topics.
-At round , a query is sampled from the task dataset.
-Players generate responses , which are pooled into .
+At round , a query is uniformly sampled from the task dataset.
+The -th player generate responses , which are then concatenated into .
 Graders then evaluate these responses, producing , which are averaged across graders to yield .
-After rounds, the cumulative evaluations are distilled into:
+After rounds, the cumulative evaluations are distilled into a confidence-weight table:
 capturing both the estimated quality and the associated uncertainty of each player.
@@
 This design calibrates the committee’s trust in different players without retraining them, enabling adaptability and efficiency.
+The confidence-weight table (index represents agent dimension, index represents task dimension)serves as a structured memory of each player’s historical reliability across subtasks (ScienceEval) or rubric dimensions (ScienceNarrative).
+Unlike fine-tuning approaches, Roundtable Policy does not alter the parameters of the underlying LLMs.
+Instead, is iteratively updated from repeated rounds of evaluation, much like how a committee gradually learns which members are most reliable in which topics.
 2.4 Inference with the Confidence-Weight Table
-Once has been learned, it guides consensus fusion at inference.
+Once has been learned, it guides consensus aggregation at the inference phase.
 Given a new query , each player generates a response .
-The collection is passed to the fusion process , which aggregates them using the pretrained :
-Here, ensures that expertise is applied where it matters most.
-In MultiTask, it emphasizes domain-specific reliability (e.g., assigning higher weights to players consistently strong in mathematics).
-In SingleTask, it captures stylistic and analytical strengths across rubric dimensions, reflecting how different voices contribute to different aspects of a proposal.
+The collection is passed to the fusion agent , which aggregates them using the pre-constructed :
 The fusion process itself is not fine-tuned; it acts as an orchestrator that conditions on to produce a consensus answer.
-This mimics how a committee chair integrates both the content and credibility of reviewers before issuing a final recommendation.
-Optionally, the AI Committee can re-evaluate to assign a quality score and confidence interval, preserving interpretability and accountability throughout the process.
+This mimics how a committee chair integrates both the content and credibility of reviewers before making a final decision.
 3 Experiments
-We design experiments to systematically evaluate the effectiveness and robustness of the proposed Roundtable Policy. Our analysis is guided by the following research questions:
-RQ1: To what extent does Roundtable Policy enhance scientific reasoning and writing performance relative to individual LLMs? (see §3.1)
-RQ2: How stable and robust is Roundtable Policy over long-run evaluation regimes, and does it consistently provide more reliable performance? (see §3.2)
-RQ3: To what degree do AI graders introduce bias, and can the AI Committee mitigate these biases to achieve greater inter-grader consistency? (see §3.3)
-3.1 Performance Improvement (RQ1)
-| Setting | Model | Material Science Dataset | Others | |||||||
+We design extensive experiments to evaluate the inference-time reasoning framework Roundtable Policy. Our analysis is guided by the following research questions:
+RQ1(Efficacy): To what extent does Roundtable Policy enhance reasoning compared with baseline models in complex scientific environments? (see §3.2)
+RQ2 (Stability & Robustness): How stable and robust is Roundtable Policy across diverse hyperparameter settings and design choices, and does it consistently provide reliable performance improvements?
+(see §3.3)
+RQ3(Multi-Agent Spectrum): Does Roundtable Policy beat other multi-agent baselines? (see §3.4)
+RQ4(Grading Rationality): To what degree do AI graders introduce bias, and can the committee setting mitigate these biases to achieve greater inter-grader consistency? (see §3.5)
+3.1 Experimental Setup
+Datasets & Evaluation Metrics.
+To simulate realistic scientific reasoning environments, we consider two complementary evaluation regimes that reflect distinct but equally important aspects of scientific work: (i)ScienceEval: cross-domain, heterogeneous problem solving (Figure 4) (ii)ScienceNarrative: coherent, narrative-driven scientific writing (Figure 3).
+These settings are together designed to capture the dual demands of scientific intelligence—answering pointwise questions and constructing structured, logically consistent arguments.
+Following is details for both settings:
+| Setting | Model | Multi-Faceted Physical Reasoning | Other Domains | |||||||
 | Mech. 1 | Mech. 2 | EE | Opt. | Thermo. | Semi. | Math | Geo. | Bio. | ||
-MultiTask |
-GPT-4o | 85.894.27 | 78.265.37 | 87.714.11 | 74.595.49 | 85.144.40 | 85.584.60 | 87.812.13 | 83.552.20 | 51.536.64 |
-| Grok-2 | 83.924.07 | 76.244.96 | 85.613.92 | 76.184.99 | 82.024.10 | 82.634.39 | 88.841.70 | 77.392.28 | 51.316.25 | |
-| Claude-3.5H | 82.074.65 | 71.206.04 | 84.724.37 | 71.495.83 | 77.745.03 | 77.955.22 | 81.192.58 | 73.682.58 | 48.526.75 | |
-| GPT-4.5p | 72.866.77 | 71.466.96 | 76.676.12 | 72.786.36 | 77.185.71 | 77.686.17 | 82.502.68 | 82.592.08 | 50.636.89 | |
-| DeepSeek-chat | 77.635.35 | 72.026.14 | 81.534.97 | 72.965.79 | 79.644.87 | 79.815.23 | 72.233.16 | 78.052.41 | 47.836.57 | |
-| Claude-3.7S | 77.433.46 | 70.514.05 | 79.243.13 | 72.393.70 | 77.413.49 | 79.493.12 | 62.342.93 | 72.561.97 | 47.765.36 | |
-| Llama-3.3 | 72.145.96 | 63.866.95 | 75.335.55 | 67.136.30 | 71.235.71 | 73.875.72 | 87.431.93 | 72.772.49 | 48.646.54 | |
-| Qwen-Plus | 80.325.01 | 72.366.03 | 82.974.71 | 73.425.66 | 81.494.70 | 81.184.93 | 86.992.08 | 76.272.48 | 51.646.57 | |
-| Gemini-2.0 | 67.267.61 | 61.618.09 | 74.966.58 | 65.737.31 | 70.726.68 | 71.606.99 | 49.774.79 | 75.172.58 | 42.117.31 | |
-| RP | 87.644.26 | 81.505.36 | 89.994.00 | 81.294.96 | 86.554.09 | 91.824.34 | 94.801.01 | 85.331.63 | 47.557.02 | |
+| ScienceEval |
+GPT-4o
+85.894.27 | 78.265.37 | 87.714.11 | 74.595.49 | 85.144.40 | 85.584.60 | 87.812.13 | 83.552.20 | 51.536.64 |
+Grok-2
+83.924.07 | 76.244.96 | 85.613.92 | 76.184.99 | 82.024.10 | 82.634.39 | 88.841.70 | 77.392.28 | 51.316.25 | |
+Claude-3.5H
+82.074.65 | 71.206.04 | 84.724.37 | 71.495.83 | 77.745.03 | 77.955.22 | 81.192.58 | 73.682.58 | 48.526.75 | |
+GPT-4.5p
+72.866.77 | 71.466.96 | 76.676.12 | 72.786.36 | 77.185.71 | 77.686.17 | 82.502.68 | 82.592.08 | 50.636.89 | |
+DeepSeek-chat
+77.635.35 | 72.026.14 | 81.534.97 | 72.965.79 | 79.644.87 | 79.815.23 | 72.233.16 | 78.052.41 | 47.836.57 | |
+Claude-3.7S
+77.433.46 | 70.514.05 | 79.243.13 | 72.393.70 | 77.413.49 | 79.493.12 | 62.342.93 | 72.561.97 | 47.765.36 | |
+Llama-3.3
+72.145.96 | 63.866.95 | 75.335.55 | 67.136.30 | 71.235.71 | 73.875.72 | 87.431.93 | 72.772.49 | 48.646.54 | |
+Qwen-Plus
+80.325.01 | 72.366.03 | 82.974.71 | 73.425.66 | 81.494.70 | 81.184.93 | 86.992.08 | 76.272.48 | 51.646.57 | |
+Gemini-2.0
+67.267.61 | 61.618.09 | 74.966.58 | 65.737.31 | 70.726.68 | 71.606.99 | 49.774.79 | 75.172.58 | 42.117.31 | |
+| Ours(RP) | 87.644.26 | 81.505.36 | 89.994.00 | 81.294.96 | 86.554.09 | 91.824.34 | 94.801.01 | 85.331.63 | 47.557.02 | |
 | Setting | Model | Background | Methodology | Impact | ||||||
 | Crea. | Sci. Rig. | Log. Coh. | Crea. | Sci. Rig. | Log. Coh. | Crea. | Sci. Rig. | Log. Coh. | ||
-SingleTask |
-GPT-4o | 70.995.01 | 74.734.68 | 82.753.76 | 75.495.23 | 74.815.03 | 79.634.31 | 72.155.19 | 70.535.25 | 79.644.19 |
-| GPT-o3 | 84.073.98 | 84.253.82 | 86.463.42 | 85.923.95 | 86.703.56 | 85.843.56 | 83.904.15 | 81.524.25 | 84.863.69 | |
-| GPT-4.1 | 76.964.62 | 81.374.16 | 86.193.51 | 81.654.67 | 83.194.19 | 84.753.85 | 78.814.74 | 77.914.67 | 83.903.88 | |
-| GPT-4.1n | 77.394.55 | 80.944.15 | 85.593.53 | 80.484.69 | 82.304.19 | 83.973.85 | 77.974.74 | 77.224.65 | 82.943.90 | |
-| GPT-o3-m | 82.654.16 | 84.783.83 | 87.263.36 | 83.154.35 | 86.863.53 | 86.443.53 | 81.504.40 | 81.824.23 | 85.153.67 | |
-| GPT-4o-m | 72.784.94 | 74.644.70 | 82.193.80 | 75.445.16 | 76.184.85 | 79.814.23 | 72.785.18 | 70.695.24 | 78.824.25 | |
-| Grok-3 | 78.505.10 | 79.104.50 | 85.503.64 | 80.964.75 | 81.264.47 | 83.583.99 | 78.814.78 | 76.344.93 | 83.044.01 | |
-| Claude-3.7S | 82.474.22 | 84.733.92 | 88.173.28 | 85.524.20 | 86.653.79 | 86.893.57 | 84.534.22 | 82.134.38 | 86.413.59 | |
-| DeepSeek-chat | 84.054.20 | 83.744.12 | 88.283.19 | 84.214.31 | 84.584.13 | 86.083.58 | 83.684.23 | 81.564.48 | 86.503.46 | |
-| RP | 90.253.95 | 89.953.78 | 91.043.09 | 90.393.91 | 91.283.35 | 91.183.20 | 90.153.95 | 89.863.95 | 90.993.26 | |
-To answer RQ1, we compare Roundtable Policy (RT) against individual LLM baselines across both MultiTask and SingleTask settings.
-Table 1 summarizes overall performance, while Figures 2(a)–2(b) visualize relative gains.
-This correspondence between player ID and model follows Table 3.
-The correspondence of subtasks of MultiTask and task IDs follows Table 4.
-Quantitative Analysis.
-In the MultiTask regime, RT delivers a 13.01% average lift, outperforming 8 out of 9 single-model baselines.
-The only exception is Task 9 (Bio.), where Qwen-plus () and GPT-4o () slightly exceed RT ().
-Beyond accuracy, RT reduces uncertainty; for instance, on Task 7 (Math) the confidence interval contracts to compared with baselines’ –.
-Figure 2(a) further shows per-player improvements.
-Gains are broadly positive on Tasks 1–8, with especially large margins on reasoning-intensive subtasks—for example, +90.5% over Gemini-2.0 on Task 7.
-Smaller improvements occur when a single model already aligns closely with ground truth, while Task 9 shows marginal or negative differences—consistent with its per-task averages.
-Appendix §D.1 confirms that weaker players (e.g., Gemini-2.0) enjoy the largest lifts, while stronger models exhibit limited headroom.
-In the SingleTask regime, which evaluates structured scientific proposal writing, RT achieves the best performance in all nine rubric cells, explaining the observed 11.04% average lift.
-Each proposal comprises three sections—Background, Methodology, and Impact—scored across Creativity, Scientific Rigor, and Logical Coherence.
-Scores cluster tightly around 90–91% with narrow confidence intervals (typically –).
-Gains concentrate in the Impact section, particularly Scientific Rigor, as shown in Figure 2(b).
-Appendix §D.2 reveals positive medians across most players, with right-skewed tails indicating that even when consensus does not outperform the strongest model, it still elevates weaker drafts significantly.
-Qualitative Analysis.
-These results illustrate how RT acts like a committee process rather than a flat average.
-In MultiTask, the confidence-weight table down-weights unreliable contributions and amplifies complementary strengths, thereby “rescuing” weaker players without diluting strong ones.
-This mirrors how, in a committee, flawed proposals can still contribute useful fragments when evaluated critically.
-The large gains on reasoning-heavy tasks (e.g., Tasks 6–7) highlight how diverse error patterns across models can be combined into substantially stronger reasoning chains.
-In SingleTask writing, weaker players show the most pronounced gains in Impact and Methodology.
-For instance, GPT-4o exhibits long right tails in Impact–Scientific Rigor and Impact–Logical Coherence, while GPT-o3 shows consistently positive medians with bounded variance.
-Even strong baselines such as DS-chat rarely degrade under RT, suggesting that consensus seldom harms high-quality drafts but frequently elevates weaker ones.
-Mechanistically, the confidence-weight table enforces rubric-aware, non-redundant fusion: systematically weak contributors are down-weighted, while demonstrated strengths are amplified.
-The pattern resembles how peer reviewers specialize—some focusing on methodology, others on impact—and the editor synthesizes these perspectives into a well-rounded final decision.
-3.2 Long-Run Stability and Robustness (RQ2)
-Setting Player ID LLM # Mean Gain (%) Std (%) p-value Significance MultiTask 001 GPT-4o 36000 6.476865 77.280369 ✓ 002 Grok-2 36000 5.254293 78.231667 ✓ 003 Claude-3.5Haiku 36000 17.251754 101.839946 ✓ 004 GPT-4.5-preview 36000 18.552776 94.463211 ✓ 005 Deepseek-Chat 36000 16.048551 96.935292 ✓ 006 Claude-3.7Sonnet 36000 13.732040 121.321577 ✓ 007 Llama-3.3 36000 24.988341 123.678543 ✓ 008 Qwen-plus 36000 12.586367 91.751670 ✓ 009 Gemini-2.0 36000 46.474125 145.071773 ✓ SingleTask 001 GPT-4o 36000 27.087781 40.303119 ✓ 002 GPT-o3 36000 13.316247 148.785675 ✓ 003 GPT-4.1 36000 15.107889 73.600957 ✓ 004 GPT-4.1-nano 36000 15.805949 28.188351 ✓ 005 GPT-o3-mini 36000 9.687830 21.760829 ✓ 006 GPT-4o-mini 36000 27.206693 42.707340 ✓ 007 Grok-3 36000 16.485277 29.024328 ✓ 008 Claude-3.7Sonnet 36000 7.761157 16.993190 ✓ 009 Deepseek-Chat 36000 7.165863 7.474061 ✓
-To answer RQ2, we test whether RT’s gains remain statistically reliable over long-run evaluations.
-Using the Wilcoxon signed-rank test (§A.2) on round-level paired differences, RT consistently outperforms all baselines.
-As shown in Table 2, across all 18 models RT achieves significant improvements at , with -values typically below .
-Quantitative Analysis.
-In the MultiTask regime, RT consistently outperforms baselines with mean percentage gains ranging from +5.3% (Grok-2) to +46.5% (Gemini-2.0).
-Even strong players such as GPT-4o (+6.5%) and Claude-3.7Sonnet (+13.7%) see reliable gains, highlighting that consensus remains beneficial even when individual agents are already competitive.
-The large improvement over Gemini-2.0 illustrates how RT effectively leverages complementary strengths to elevate weaker contributors, rather than letting them dilute overall performance.
-In the SingleTask regime, RT delivers significant lifts across all nine baselines, with mean improvements from +7.2% (Deepseek-Chat) to +27.2% (GPT-4o-mini).
-Notably, gains are especially pronounced for smaller or lighter-weight models (e.g., GPT-4o-mini, GPT-4o), suggesting that RT acts as a stabilizing mechanism that brings their drafts closer to the quality of stronger models.
-Even against advanced systems like GPT-o3 and Claude-3.7Sonnet, improvements remain consistently positive, confirming robustness across a wide range of capability levels.
-Qualitative Analysis.
-Beyond statistical significance, the long-run results highlight important qualitative patterns.
-In the MultiTask regime, RT’s robustness stems from its ability to repeatedly exploit complementary error profiles: weaker players such as Gemini-2.0 or Llama-3.3 often fail in different ways, but when their contributions are weighted through the confidence table, the ensemble consistently extracts correct elements while suppressing noise.
-This is reflected in the large variance reported in Table 2—although individual rounds fluctuate, the aggregated effect over thousands of rounds converges to stable gains.
-In committee terms, even if certain members are unreliable in a given discussion, their occasional correct insights are preserved while systematic mistakes are filtered out, ensuring dependable group performance.
-In the SingleTask regime, stability is most visible at the rubric level.
-Analytical dimensions such as Impact–Scientific Rigor show the largest sustained gains, whereas stylistic traits (e.g., Creativity or Logical Coherence) improve more modestly.
-This asymmetry suggests that RT is especially effective in consolidating rigorous, reasoning-heavy contributions—akin to how editors place greater weight on technical soundness than on stylistic flair when synthesizing reviewer reports.
-Importantly, the consistency of improvements across tens of thousands of rounds demonstrates that RT does not rely on one-off synergies, but instead behaves like a well-tuned review board: repeatedly integrating diverse perspectives into a balanced and accountable consensus.
-3.3 Bias and Consistency of AI Graders (RQ3)
-To address RQ3, we examine grader bias and inter-grader consistency through two complementary lenses: (i) the distribution of ranks assigned to RT by different graders, and (ii) pairwise agreement measured by Kendall’s Tau (see §A.3).
-Aggregated results are shown in Figure 3, with task- and dimension-level breakdowns provided in Figures 14 and 13 in §D.3.
-Quantitative Analysis.
-Figure 3 highlights a clear distinction between MultiTask and SingleTask regimes.
-In MultiTask, the aggregated rank distribution is steep and concentrated in the top three positions, with RT frequently ranked first.
-Correspondingly, Kendall’s Tau values are consistently high (–), reflecting strong alignment among graders.
-Task-level inspection (Figures 14, 13) confirms this pattern: Tasks 001–006 exhibit overlapping distributions, Tasks 007–008 achieve near-perfect consensus (), while Task 009 shows the weakest agreement (–) as some graders disperse RT’s rank across mid-range positions.
-In SingleTask, however, the aggregated rank distribution is flatter and more dispersed, and Kendall’s Tau values drop markedly (– on average).
-Creativity-related dimensions (e.g., Background–Creativity, Methodology–Creativity) place RT near the top but with substantial grader variability.
-Logical coherence dimensions show the lowest alignment, with Tau values as low as –, indicating inconsistent judgments.
-By contrast, Impact–Scientific Rigor yields the strongest agreement, with RT concentrated near rank 1 and Tau values of –.
-Together, these results demonstrate that graders achieve high consensus in objective, structured tasks but diverge significantly when assessing subjective, open-ended proposals.
-Qualitative Analysis.
-These results reveal that bias is minimal in structured MultiTask settings, where evaluation criteria are well defined, but more pronounced in open-ended SingleTask writing, where subjective interpretation shapes grading.
-This mirrors human peer review: reviewers evaluating technical correctness tend to agree, while judgments about creativity or style diverge.
-Crucially, the AI Committee mitigates such variability by aggregating across graders, reducing the impact of individual biases.
-In practice, this means RT functions like an editorial board: even when reviewers disagree, their collective judgment yields a more balanced and reliable outcome.
-4 Ablation Study
-To disentangle the contributions of our design choices and validate their scientific relevance, we conduct two ablation studies.
-These analyses illustrate how Roundtable Policy strengthens structured reasoning and improves narrative quality by highlighting what is gained—and what is lost—when key components are removed.
-Policy Ablation.
-We first compare RT with a Best-Single-Model Policy (BSMP), which simply selects the response from the single model with the highest confidence weight rather than synthesizing across multiple players.
-Conceptually, BSMP resembles a committee that crowns one “top performer” and ignores all other voices.
-The results in Figure 4 make clear why this approach falls short: across both MultiTask and SingleTask settings, RT consistently outperforms BSMP.
-In MultiTask, the advantages are particularly evident in reasoning-heavy domains such as mathematics and mechanics, where drawing on complementary perspectives yields large gains.
-In SingleTask proposal writing, consensus fusion not only raises average scores across all rubric dimensions but also reduces variance, producing outputs that are both stronger and more consistent.
-These findings highlight that scientific progress rarely emerges from a single dominant voice; it is the synthesis of multiple viewpoints that produces reliable reasoning and coherent narrative.
-Rounds Ablation.
-We next examine the impact of the number of training rounds used to construct the confidence-weight table.
-Here, the table plays the role of institutional memory—accumulating evidence about which players are dependable under different conditions.
-As shown in Figure 5, the effect of this memory differs by task type.
-In MultiTask, where benchmarks span diverse scientific domains, performance steadily improves as more rounds are incorporated and stabilizes only after hundreds of evaluations.
-By contrast, SingleTask writing converges much more quickly: fewer than 100 rounds already suffice to surpass baselines, with further rounds offering only modest additional benefit.
-This divergence reflects a natural distinction: heterogeneous domains require extensive calibration to build trust across disciplines, whereas rubric-based writing tasks benefit from faster stabilization since evaluative criteria are more homogeneous.
+| ScienceNarrative |
+GPT-4o
+70.995.01 | 74.734.68 | 82.753.76 | 75.495.23 | 74.815.03 | 79.634.31 | 72.155.19 | 70.535.25 | 79.644.19 |
+GPT-o3
+84.073.98 | 84.253.82 | 86.463.42 | 85.923.95 | 86.703.56 | 85.843.56 | 83.904.15 | 81.524.25 | 84.863.69 | |
+GPT-4.1
+76.964.62 | 81.374.16 | 86.193.51 | 81.654.67 | 83.194.19 | 84.753.85 | 78.814.74 | 77.914.67 | 83.903.88 | |
+GPT-4.1n
+77.394.55 | 80.944.15 | 85.593.53 | 80.484.69 | 82.304.19 | 83.973.85 | 77.974.74 | 77.224.65 | 82.943.90 | |
+GPT-o3-m
+82.654.16 | 84.783.83 | 87.263.36 | 83.154.35 | 86.863.53 | 86.443.53 | 81.504.40 | 81.824.23 | 85.153.67 | |
+GPT-4o-m
+72.784.94 | 74.644.70 | 82.193.80 | 75.445.16 | 76.184.85 | 79.814.23 | 72.785.18 | 70.695.24 | 78.824.25 | |
+Grok-3
+78.505.10 | 79.104.50 | 85.503.64 | 80.964.75 | 81.264.47 | 83.583.99 | 78.814.78 | 76.344.93 | 83.044.01 | |
+Claude-3.7S
+82.474.22 | 84.733.92 | 88.173.28 | 85.524.20 | 86.653.79 | 86.893.57 | 84.534.22 | 82.134.38 | 86.413.59 | |
+DeepSeek-chat
+84.054.20 | 83.744.12 | 88.283.19 | 84.214.31 | 84.584.13 | 86.083.58 | 83.684.23 | 81.564.48 | 86.503.46 | |
+| Ours(RP) | 90.253.95 | 89.953.78 | 91.043.09 | 90.393.91 | 91.283.35 | 91.183.20 | 90.153.95 | 89.863.95 | 90.993.26 | |
+ScienceEval: Agents are simultaneously queried with nine domain-specific questions spanning geoscience (Deng et al., 2024), biology (Jin et al., 2019), mathematics (Cobbe et al., 2021), and multi-faceted physical science(Appendix §C), containing six sub-categories.
+The tasks include multiple-choice, fill-in-the-blank, and short-answer formats.
+This setting resembles a comprehensive written examination covering diverse scientific disciplines, where success requires broad and balanced reasoning capabilities rather than specialization in a single domain.
+ScienceNarrative: Agents are required to generate structured scientific proposals following a standardized three-paragraph template consisting of Background, Methodology, and Impact (§C).
+This regime emphasizes the narrative dimension of scientific reasoning: beyond producing isolated facts, agents must organize ideas into a coherent, logically consistent, and persuasive scientific narrative.
+Baselines.
+We evaluate Roundtable Policy against both single-agent and multi-agent baselines.
+As single-agent baselines, we adopt a diverse set of state-of-the-art LLMs following the notation in Table 4.
+For multi-agent collaboration, we consider representative frameworks: Majority Vote (Wang et al., 2023) aggregates agents by simple voting and serves as a canonical non-weighted baseline.
+Weighted Vote (Taubenfeld et al., 2025) extends majority voting by assigning confidence-based weights to individual agents.
+Debate (Du et al., 2024) enables agents to iteratively revise their answers through argument exchange.
+Debate w Judge (Liang et al., 2024) introduces an explicit adjudicator to select the final answer after debate, while Debate w Confidence (Chen et al., 2024) further incorporates confidence signals to guide resolution.
+Additionally, Mixture-of-Agents (Wang et al., 2025) constructs an agentic network to combine multiple agent outputs adaptively.
+3.2 Performance Improvement (RQ1)
+To evaluate the efficacy, we compare Roundtable Policy (RP) against single-LLM baselines across both the ScienceEval and ScienceNarrative.
+Table 1 summarizes overall performance, while Figures 13 and 15 visualize fine-grained percentage improvements.
+Additional per-player and per-task analyses are provided in Appendix §D.1 and §D.2.
+Based on these results, we draw the observations:
+Obs 1: RP consistently improves performance across heterogeneous scientific tasks, with particularly strong gains on reasoning-intensive problems, while maintaining stable and reliable outcomes.
+In the ScienceEval, RP achieves a 13.01% average improvement over single-model baselines.
+Performance gains are broadly positive on Tasks 1–8 and are especially pronounced on reasoning-heavy subtasks (multi-faceted physical reasoning).
+Beyond mean accuracy, RP also reduces predictive uncertainty; for example, on Task 7 (Math), RP narrows the confidence interval to , compared with – for single-model baselines.
+In contrast, Task 9 (Biology) exhibits marginal or negative differences, likely due to high intrinsic uncertainty and substantial disagreement among individual models, which limits the effectiveness of consensus-based aggregation.
+These results indicate that RP is most effective when model errors are diverse yet structured.
+Obs 2: RP yields robust and uniform improvements in structured scientific writing, substantially elevating weaker drafts without degrading strong ones.
+In the ScienceNarrative, RP achieves superior performance across all nine rubric dimensions, with an 11.04% average improvement.
+Gains concentrate in the Impact section—particularly Scientific Rigor—while overall scores remain tightly clustered around 90–91% with narrow confidence intervals.
+Per-player analyses reveal right-skewed improvement distributions, indicating that RP consistently lifts weaker proposals while preserving the quality of already strong drafts.
+3.3 Stability and Robustness (RQ2)
+Setting
+Model
+Avg. Gain (%)
+Std (%)
+p-value
+Significance
+ScienceEval
+-4o
+6.476865
+77.280369
+-2
+5.254293
+78.231667
+-3.5H
+17.251754
+101.839946
+-4.5p
+18.552776
+94.463211
+-Chat
+16.048551
+96.935292
+-3.7S
+13.732040
+121.321577
+-3.3
+24.988341
+123.678543
+-plus
+12.586367
+91.751670
+-2.0
+46.474125
+145.071773
+ScienceNarrative
+-4o
+27.087781
+40.303119
+-o3
+13.316247
+148.785675
+-4.1
+15.107889
+73.600957
+-4.1-nano
+15.805949
+28.188351
+-o3-mini
+9.687830
+21.760829
+-4o-mini
+27.206693
+42.707340
+-3
+16.485277
+29.024328
+-3.7S
+7.761157
+16.993190
+-Chat
+7.165863
+7.474061
+We examine whether the performance gains are statistically reliable and robust under long-run interactions and diverse design choices.
+Specifically, we conduct Wilcoxon signed-rank tests to assess statistical significance and perform systematic ablations to evaluate structural robustness (Figure 5).
+Obs 3: RP provides statistically reliable and long-run stable performance improvements.
+As shown in Table 2, RP significantly outperforms all agents at , with -values typically below .
+Despite substantial per-round variance, the aggregated gains consistently converge to stable improvements over of thousands of evaluations.
+This phenomenon indicates that RP’s advantages are not driven by isolated cases or short-lived fluctuations, but reflect persistent distribution-level improvements.
+Fine-grained Wilcoxon results are reported in Tables 11, 12, 9 and 10.
+Obs 4: RP’s stability is structurally grounded and manifests consistently across multiple design parameters.
+Figure 5 shows that RP remains robust under systematic ablations of various components.
+Scaling the number of player agents leads to steady performance improvements that gradually saturate.
+In practice, the number of players is constrained by the maximum context window of the aggregator, as all proposals are concatenated for aggregation; we therefore evaluate scalability within this realistic operating regime.
+Second, ablating the aggregator from strong to weak results in bounded performance degradation rather than drastic collapse, indicating that RP does not rely on a single highly capable aggregator.
+Additionally, increasing the number of pretraining rounds111historical rounds to distill the confidence-weight table yields consistent gains, indicating that more accurate encoding of historical reliability leads to better downstream performance.
+Finally, incorporating uncertainty-aware scoring outperforms vanilla score aggregation, reflecting that RP’s reasoning improvements are not driven by fragile or ad hoc scoring heuristics.
+3.4 Multi-Agent Baselines (RQ3)
+| Setting | Method | Multi-Faceted Physical Reasoning | Other Domains | |||||||
+| Mech. 1 | Mech. 2 | EE | Opt. | Thermo. | Semi. | Math | Geo. | Bio. | ||
+| ScienceEval | Majority Vote (Wang et al., 2023) | 82.484.51 | 74.025.63 | 85.014.25 | 73.655.48 | 81.014.56 | 81.684.77 | 85.932.12 | 79.032.27 | 50.886.54 |
+| Weighted Vote (Taubenfeld et al., 2025) | 81.564.58 | 74.465.54 | 84.244.29 | 73.135.45 | 81.764.49 | 81.164.77 | 83.032.27 | 76.562.31 | 49.286.48 | |
+| Debate (Du et al., 2024) | 85.684.12 | 80.344.28 | 88.014.02 | 77.253.90 | 84.973.30 | 89.174.09 | 92.231.05 | 84.051.90 | 51.686.25 | |
+| Debate w Judge (Liang et al., 2024) | 86.934.28 | 80.364.20 | 88.424.10 | 76.254.05 | 85.263.77 | 89.434.25 | 92.041.25 | 83.651.97 | 51.236.07 | |
+| Debate w Confidence (Chen et al., 2024) | 86.243.50 | 79.613.69 | 88.153.55 | 79.393.73 | 86.273.02 | 89.953.45 | 93.181.55 | 83.921.30 | 49.174.75 | |
+| Mixture-of-Agents (Wang et al., 2025) | 87.533.75 | 81.023.74 | 89.743.47 | 80.683.23 | 87.293.05 | 91.253.59 | 93.051.42 | 85.281.25 | 53.505.31 | |
+| Roundtable Policy | 87.644.26 | 81.505.36 | 89.994.00 | 81.294.96 | 86.554.09 | 91.824.34 | 94.801.01 | 85.331.63 | 47.557.02 | |
+| Setting | Model | Background | Methodology | Impact | ||||||
+| Crea. | Sci. Rig. | Log. Coh. | Crea. | Sci. Rig. | Log. Coh. | Crea. | Sci. Rig. | Log. Coh. | ||
+| ScienceNarrative | Majority Vote (Wang et al., 2023) | 81.374.30 | 82.994.07 | 86.963.41 | 83.464.37 | 84.643.98 | 85.433.70 | 81.934.41 | 80.114.51 | 84.903.73 |
+| Weighted Vote (Taubenfeld et al., 2025) | 81.474.30 | 83.224.05 | 87.263.36 | 83.744.36 | 84.903.97 | 85.743.68 | 82.374.38 | 80.354.51 | 85.223.70 | |
+| Debate (Du et al., 2024) | 83.164.88 | 84.995.25 | 88.253.64 | 86.145.25 | 87.203.50 | 86.923.25 | 85.114.02 | 83.654.00 | 85.963.75 | |
+| Debate w Judge (Liang et al., 2024) | 86.753.75 | 85.504.25 | 86.044.17 | 87.053.50 | 88.963.25 | 88.813.10 | 87.254.06 | 83.503.85 | 85.413.71 | |
+| Debate w Confidence (Chen et al., 2024) | 86.473.20 | 89.253.34 | 90.022.95 | 86.993.13 | 90.253.25 | 90.702.75 | 87.083.50 | 88.633.31 | 89.822.95 | |
+| Mixture-of-Agents (Wang et al., 2025) | 89.983.90 | 86.174.47 | 89.503.68 | 90.253.65 | 88.923.25 | 90.153.57 | 89.754.25 | 88.603.93 | 89.223.10 | |
+| Roundtable Policy | 90.253.95 | 89.953.78 | 91.043.09 | 90.393.91 | 91.283.35 | 91.183.20 | 90.153.95 | 89.863.95 | 90.993.26 | |
+Furthermore, we compare Roundtable Policy with representative multi-agent frameworks (implementation details at Appendix §E), including voting-based aggregation (Wang et al., 2023; Taubenfeld et al., 2025), debate-style interaction (Du et al., 2024; Liang et al., 2024; Chen et al., 2024), and agentic-network-style aggregation (Wang et al., 2025).
+Beyond raw performance, we analyze their behaviors through the lens of how agent influence is assigned, updated, and stabilized during aggregation.
+Obs 5: RP achieves consistently strong and balanced performance in complex reasoning environments.
+In the ScienceEval, RP attains the best or second-best results on 8 out of 9 task dimensions, with particularly strong performance on reasoning-intensive domains such as multi-faceted physical reasoning and Math.
+This behavior contrasts with voting-based methods, whose uniform and memoryless aggregation is prone to majority bias when partial but confident opinions dominate.
+While interaction-heavy baselines such as Debate and Debate w Confidence can achieve competitive or even superior performance on individual tasks, their gains are highly task-dependent and exhibit noticeable instability, reflecting sensitivity to transient interaction dynamics.
+The agentic-network-style method (Mixture-of-Agents) also demonstrates strong performance, which we attribute to the rich architectural expressivity: layered agents integrate diverse perspectives, analogous to a highly expressive super-neuron network.
+However, this expressivity comes with structural redundancy, and the aggregation mechanism remains largely static, limiting its ability to adaptively recalibrate agent influence under diverse environments.
+Obs 6: RP demonstrates superior generality in complex scientific writing.
+In the ScienceNarrative, RP achieves best performance across all nine rubrics.
+Debate-based methods excel in dimensions emphasizing argumentative clarity or coherence, but their performance degrades when tasks require creativity.
+Debate with Confidence partially mitigates this issue by incorporating confidence cues during interaction; however, such cues remain interaction-conditioned and often lead to over-neutralization.
+As shown in Table 3, competing methods exhibit strengths in specific categories, whereas RP consistently maintains balanced performance across all rubrics.
+3.5 Bias and Consistency (RQ4)
+We analyze grader bias and inter-grader consistency using (i) rank distributions assigned to RP and (ii) pairwise agreement measured by Kendall’s Tau (Appendix §A.3).
+Aggregate results are shown in Figure 7, with detailed breakdowns in Appendix §D.3.
+Obs 7: AI graders are consistent on structured tasks and more variable on open-ended evaluations, while committee aggregation mitigates individual bias.
+In ScienceEval, graders exhibit concentrated rank distributions and high Kendall’s Tau values, indicating strong agreement on objective tasks.
+In ScienceNarrative, inter-grader agreement decreases, particularly for subjective dimensions such as creativity and logical coherence.
+Nevertheless, aggregating across multiple graders substantially reduces individual bias, yielding balanced assessments.
 Summary.
-These ablations underscore two core insights: First, weighted consensus is fundamentally more powerful than relying on a single “best” model, as robust science arises from collective deliberation rather than solitary voices.
-Second, the effectiveness of the confidence-weight table depends on task structure: structured MultiTask settings demand long-run calibration, while open-ended SingleTask writing achieves stability with much less historical data.
-Together, these findings reinforce that consensus-driven fusion is not only effective but also principled—mirroring how real scientific communities build reliable knowledge through collective judgment.
+By decoupling influence assignment from transient interaction dynamics and aligning long-term fine-grained measurement, RP enables stable context-aware aggregation.
+In addition to increasingly complex interaction protocols, our observations find that multi-agent systems also benefit from mechanisms that explicitly model, update, and stabilize agent reliability across downstream settings.
+Future inference-time reasoning frameworks for multi-agent systems may increasingly view collaboration as a problem of reliability modeling and consensus formation, rather than relying solely on more elaborate interaction dynamics.
 4 Related Work
-LLMs for Scientific Discovery.
-Large language models (LLMs) have rapidly expanded the scope of data-driven science, from AlphaFold’s breakthroughs in structural biology (Jumper et al., 2021) to recent applications across mathematics (Yue et al., 2024), chemistry (Yu et al., 2024), biology (Labrak et al., 2024), and geography (Li et al., 2023).
-Their strength lies in structured reasoning and seamless integration with external tools, enabling automated workflows (Cao, 2017).
-Yet, complex multi-step reasoning often requires specialized prompting, tool-augmented inference, or fine-tuning, motivating the rise of agentic systems that combine reasoning with external solvers and databases (Boiko et al., 2023; Lu et al., 2024).
-Reasoning and Factuality in Language Models.
-Improving reasoning and factuality in LLMs has been studied extensively.
-Reasoning can be enhanced through prompting methods such as scratchpads (Nye & et al., 2021), chain-of-thought demonstrations (Wei et al., 2022; Kojima et al., 2022; Reynolds & McDonell, 2021), verification and self-consistency (Cobbe et al., 2021; Wang et al., 2022), and reflection (Shinn et al., 2023; Madaan & et al., 2023), as well as fine-tuning approaches (Rajani et al., 2019; Zelikman et al., 2022; Lewkowycz et al., 2022).
-Factuality, in turn, has been strengthened through RLHF (Christiano et al., 2017; Ziegler & et al., 2019; Ouyang et al., 2022), dataset curation (Lee & et al., 2022), retrieval-based augmentation (Guu et al., 2020), and likelihood-based calibration (Kadavath & et al., 2022).
-Multi-Agent and Ensemble Approaches.
-Beyond single-model prompting, researchers have explored ensemble and multi-agent strategies to enhance reliability.
-Early work shows that self-consistency (Wang et al., 2022) and majority voting (Li et al., 2022; Cobbe et al., 2021; Thoppilan et al., 2022) can mitigate errors by aggregating diverse reasoning chains.
-Building on this, multi-agent debate frameworks (Irving et al., 2018; Du & et al., 2023; Madaan & et al., 2023) and collaborative dialogue methods (Chan & et al., 2023; Wang & et al., 2023) let models iteratively critique and refine one another’s outputs.
-Related lines of work also investigate compositional generation, combining outputs from multiple models (Du et al., 2020; Liu & et al., 2022; Alayrac & et al., 2022; Zeng & et al., 2022).
-While these methods demonstrate the value of collective intelligence, they often rely on flat aggregation or emergent convergence.
-In contrast, our approach introduces a structured consensus mechanism that explicitly weights models by demonstrated reliability, producing interpretable and auditable signals of trust.
-Benchmarks for LLM-Based Agents.
-Evaluating agentic LLMs has inspired benchmarks ranging from synthetic tool-use tasks (e.g., TaskBench (Shen et al., 2024)) to domain-specific challenges such as SWE-Bench (Jimenez et al., 2024), BioCoder-Py (Tang et al., 2024), and MLAgentBench (Huang et al., 2024).
-More recently, datasets sourced from scientific literature (Majumder et al., 2024) aim to capture authentic workflows.
-Our work complements these efforts by introducing a domain-authentic benchmark in materials science that stresses both mechanistic understanding and property prediction, thereby testing agents under both interpretative and predictive dimensions.
-Machine Learning and Agents for Materials Science.
-Machine learning has already transformed materials science by enabling scalable property prediction, compound discovery, and structure–property analysis (Jain et al., 2013; Butler et al., 2018), with advances in reinforcement learning for inverse design (Segler et al., 2018), generative modeling (Granda et al., 2018), and multi-modal integration of simulations and experiments (Pyzer-Knapp et al., 2022).
-These developments reduce experimental costs and provide interpretable insights, but LLM-based agents remain relatively underexplored in this domain.
-The inherent complexity of materials science—multiscale interactions, vast compositional spaces, and intricate synthesis pipelines—calls for agent frameworks that unify predictive modeling with experimental feedback.
-Early systems in autonomous discovery (Gregoire et al., 2023; Maurizi et al., 2024) show promise, but important gaps remain in linking reasoning, planning, and uncertainty quantification.
-Our work positions materials science as a natural testbed for stress-testing multi-agent reasoning frameworks such as Roundtable Policy, bridging structured evaluation and open-ended scientific narrative.
-6 Conclusion & Future Directions
-We presented the Roundtable Policy, a inference-time reasoning framework that fuses outputs from multiple LLMs through confidence-weighted aggregation without finetuning base models.
-Unlike single-model prompting or naive ensemble methods such as majority voting, our approach does not treat all models equally nor rely on opaque emergent convergence, as in debate-style methods.
-Instead, it introduces a principled mechanism that explicitly encodes historical reliability and uncertainty, producing consensus answers that are not only more accurate but also transparent and auditable.
-Across structured MultiTask benchmarks and open-ended SingleTask proposal writing, we showed that Roundtable Policy consistently improves factual reliability, reasoning consistency, and narrative coherence, highlighting its value as a trustworthy tool for scientific reasoning.
-Looking forward, several avenues can extend this framework.
-First, adaptive arbitrator strategies could allow the confidence-weight table to evolve dynamically with task complexity, reallocating trust in real time.
-Second, coupling the fusion process with reinforcement learning and human or experimental feedback may enable autonomous optimization of weight assignments through trial and error.
-Third, incorporating richer feedback signals—such as calibrated uncertainty, tool-augmented checks, or domain-specific rubrics—could further strengthen robustness in complex scientific workflows.
-More broadly, the principles underlying Roundtable Policy generalize beyond the settings studied here.
-By tailoring consensus mechanisms to domain-specific evaluation criteria, the framework can support applications ranging from drug discovery and computational chemistry to astrophysics and other data-intensive sciences.
-In this sense, Roundtable Policy is not merely another ensemble method, but a scalable and interpretable paradigm for collective intelligence in LLMs—one that advances the rigor, reliability, and accountability of AI-assisted scientific discovery.
-Ethics Statement
-Our method significantly enhances the performance of sequential model editing, making it invaluable for updating and managing knowledge in real-world applications.
-While the ability to directly modify stored knowledge introduces potential risks, such as the introduction of false or harmful information, we strongly urge researchers to implement strict validation and oversight to ensure the ethical use of these techniques.
-Nevertheless, the original goal of model editing is positive, aiming to facilitate efficient updates of large models in the future.
-Therefore, we encourage researchers to leverage this technology responsibly and with care.
-Reproducibility
-To ensure the reproducibility of our findings, detailed implementation instructions for Roundtable Policy can be found in Appendix.
-Additionally, the source code is publicly available at https://ai4scidisc.github.io/roundtable/.
-These measures are intended to facilitate the verification and replication of our results by other researchers in the field.
-Acknowledement
-We gratefully acknowledge the MIT SuperCloud and Texas Advanced Computing Center (TACC) at The University of Texas at Austin for providing computational resources that have contributed to the research results reported within this paper.
-JL acknowledges support by the Defense Advanced Research Projects Agency (DARPA) under Agreement No.
-HR00112490369.
-The content is solely the responsibility of the authors and does not necessarily represent the official views of the funding entities.
+Multi-Agent Systems.
+Early approaches show that majority (Wang et al., 2023) and confidence-weighted schemes (Taubenfeld et al., 2025) can improve reliability by combining diverse reasoning trajectories.
+Besides, Du et al.
+(2024) present multi-agent debate, while Liang et al.
+(2024) introduce asymmetric debater–judge roles in debate.
+Further collaborative paradigm emphasize fine-grained interaction protocols (Chan et al., 2024; Zhang et al., 2024; Wang et al., 2025).
+Our work provides a complementary perspective by studying how reliability-aware consensus can be used during inference, compatible with other collaboration strategies.
+LLMs for Science.
+Large foundation models have driven broad scientific advances, from protein structure prediction (Senior et al., 2020; Jumper et al., 2021; Baek et al., 2021; Lin et al., 2023; Elnaggar et al., 2021) to domain-specialized tasks spanning math, chemistry, biology, and geoscience (Romera-Paredes et al., 2024; Trinh et al., 2024; Yue et al., 2024; Yu et al., 2024; Labrak et al., 2024; Li et al., 2023).
+Our work complements increasingly complex agentic workflows that couple LLMs with external tools and experimental pipelines (Yao et al., 2023; Boiko et al., 2023; Lu et al., 2024; Yamada et al., 2025) by focusing on end-to-end inference-time reasoning and aggregation.
+Benchmarks for Scientific Reasoning.
+Scientific reasoning has been widely benchmarked.
+Math GSM8K (Cobbe et al., 2021) focus on step-by-step numerical reasoning, biomedical PubMedQA (Jin et al., 2019) assess Q&A grounded in literature, GPQA and AlphaGeometry (Trinh et al., 2024) introduce problem-solving in specialized settings.
+Multi-domain benchmarks such as MMLU (Hendrycks et al., 2021) and BIG-Bench (bench authors, 2023) evaluate general reasoning across a broad range of technical subjects, ARC (Clark et al., 2018), GSM-Symbolic (Mirzadeh et al., 2025), and GPQA-Diamond (Rein et al., 2024) further emphasize compositional reasoning in diverse settings.
+Our work extends these efforts by integrating heterogeneous, long-context scientific reasoning, providing explicit dimensions for assessing expertise in multi-agent systems.
+5 Conclusion
+We presented Roundtable Policy, an inference-time reasoning framework for multi-agent systems, and demonstrated its efficacy through extensive experiments.
+More broadly, this work suggests that multi-agent collaboration for complex tasks may benefit from explicitly modeling and stabilizing agent reliability and consensus, motivating principled inference-time strategies, and evaluation frameworks that better reflect the characteristics of downstream tasks.
+Impact Statement
+Roundtable Policy can boost performance in diverse downstream tasks using multi-agent systems.
 Appendix A Notation and Definitions
@@
 : the query space from which tasks are sampled.
+: a general dataset.
 : dataset consisting of subtasks with known ground truths (e.g., multiple-choice, fill-in-the-blank, short-answer).
 : dataset of structured scientific proposal prompts without ground truth, each consisting of three required paragraphs.
-: the number of player agents (pretrained LLMs generating candidate responses). In our experiments .
+: the number of player agents (proposer agents, pretrained LLMs generating candidate responses). In our experiments .
 : the number of grader agents forming the AI Committee (evaluators of player responses). In our experiments .
@@
 : the -th grader agent, producing an evaluation score and uncertainty tuple .
-: the fusion agent, which synthesizes a final consensus answer from all player responses using the confidence-weight table.
+: the fusion agent (aggregator agent), which synthesizes a final consensus answer from all player responses using the confidence-weight table.
 : the query (or query batch) at round .
-: the corresponding reference answers (available only in MultiTask).
+: the corresponding reference answers (available only in ScienceEval).
 : the response generated by player in round .
@@
 : the confidence-weight table, a matrix storing trust weights for each player and subtask/rubric dimension.
-: the fused consensus response in round .
-| Setting | Player ID | Abbreviation | Full Name | Citation |
-MultiTask |
-001 | GPT-4o | GPT-4o | OpenAI (2024b) |
-| 002 | Grok-2 | Grok-2 | xAI (2024) | |
-| 003 | Claude-3.5H | Claude-3.5 Haiku | Anthropic (2024) | |
-| 004 | GPT-4.5p | GPT-4.5-preview | OpenAI (2024a) | |
-| 005 | DS-chat | DeepSeek-Chat | AI (2024) | |
-| 006 | Claude-3.7S | Claude-3.7 Sonnet | Anthropic (2025) | |
-| 007 | Llama-3.3 | Llama-3.3 | AI (2025) | |
-| 008 | Qwen+ | Qwen-plus | Team (2024) | |
-| 009 | Gemini-2.0 | Gemini-2.0 | DeepMind (2024) | |
-SingleTask |
-001 | GPT-4o | GPT-4o | OpenAI (2024b) |
-| 002 | GPT-o3 | GPT-o3 | OpenAI (2024d) | |
-| 003 | GPT-4.1 | GPT-4.1 | OpenAI (2023a) | |
-| 004 | GPT-4.1n | GPT-4.1-nano | OpenAI (2023b) | |
-| 005 | GPT-o3-m | GPT-o3-mini | OpenAI (2024e) | |
-| 006 | GPT-4o-m | GPT-4o-mini | OpenAI (2024c) | |
-| 007 | Grok-3 | Grok-3 | xAI (2025) | |
-| 008 | Claude-3.7S | Claude-3.7 Sonnet | Anthropic (2025) | |
-| 009 | DS-chat | DeepSeek-Chat | AI (2024) | |
-| Setting | Grader ID | Abbreviation | Full Name | Citation |
-MultiTask |
-001 | GPT-4o | GPT-4o | OpenAI (2024b) |
-| 002 | GPT-4.1 | GPT-4.1 | OpenAI (2023a) | |
-| 003 | GPT-o3 | GPT-o3 | OpenAI (2024d) | |
-| 004 | GPT-o1 | GPT-o1 | OpenAI (2023c) | |
-SingleTask |
-001 | GPT-4o | GPT-4o | OpenAI (2024b) |
-| 002 | GPT-4.1 | GPT-4.1 | OpenAI (2023a) | |
-| 003 | GPT-o3 | GPT-o3 | OpenAI (2024d) | |
-| 004 | GPT-o1 | GPT-o1 | OpenAI (2023c) |
+: the fused consensus response in round , using a pre-constructed confidence-weight table.
+| Setting | Player ID | Abbreviation | Full Name |
+| SciEval | 001 | GPT-4o | GPT-4o |
+| 002 | Grok-2 | Grok-2 | |
+| 003 | Claude-3.5H | Claude-3.5 Haiku | |
+| 004 | GPT-4.5p | GPT-4.5-preview | |
+| 005 | DS-chat | DeepSeek-Chat | |
+| 006 | Claude-3.7S | Claude-3.7 Sonnet | |
+| 007 | Llama-3.3 | Llama-3.3 | |
+| 008 | Qwen+ | Qwen-plus | |
+| 009 | Gemini-2.0 | Gemini-2.0 | |
+| SciNarrative | 001 | GPT-4o | GPT-4o |
+| 002 | GPT-o3 | GPT-o3 | |
+| 003 | GPT-4.1 | GPT-4.1 | |
+| 004 | GPT-4.1n | GPT-4.1-nano | |
+| 005 | GPT-o3-m | GPT-o3-mini | |
+| 006 | GPT-4o-m | GPT-4o-mini | |
+| 007 | Grok-3 | Grok-3 | |
+| 008 | Claude-3.7S | Claude-3.7 Sonnet | |
+| 009 | DS-chat | DeepSeek-Chat | |
+| Setting | Grader ID | Abbreviation | Full Name |
+| SciEval | 001 | GPT-4o | GPT-4o |
+| 002 | GPT-4.1 | GPT-4.1 | |
+| 003 | GPT-o3 | GPT-o3 | |
+| 004 | GPT-o1 | GPT-o1 | |
+| SciNarrative | 001 | GPT-4o | GPT-4o |
+| 002 | GPT-4.1 | GPT-4.1 | |
+| 003 | GPT-o3 | GPT-o3 | |
+| 004 | GPT-o1 | GPT-o1 | |
+| Setting | Aggregator Type | Full Name | |
+| SciEval | Strong | GPT-4o | |
+| Intermediate-1 | Llama-3.3 | ||
+| Intermediate-2 | Qwen-Plus | ||
+| Weak | Gemini-2.0 | ||
+| SciNarrative | Strong | Claude-3.7S | |
+| Intermediate-1 | Grok-3 | ||
+| Intermediate-2 | GPT-4.1 | ||
+| Weak | GPT-4o | ||
 A.2 Formal Hypothesis Testing Framework for Wilcoxon Signed-Rank Significance Validity
-To rigorously assess whether the Roundtable Policy (RT) provides statistically significant improvements over individual LLM baselines, we apply the non-parametric Wilcoxon signed-rank test to paired, round-level performance data.
+To rigorously assess whether the Roundtable Policy (RP) provides statistically significant improvements over individual LLM baselines, we apply the non-parametric Wilcoxon signed-rank test to paired, round-level performance data.
 This test evaluates whether the median of the differences between two matched samples significantly exceeds zero, under the null hypothesis of no improvement.
@@
 We define:
-: score obtained by RT in round ,
+: score obtained by RP in round ,
 : score obtained by player in the same round,
-: relative gain of RT over player in round (as percentage).
+: relative gain of RP over player in round (as percentage).
 Hypothesis Test.
@@
 This approximation is used in our evaluation whenever exceeds the small-sample regime (), ensuring computational efficiency without sacrificing statistical validity.
-A.2.1 Application to MultiTask Evaluation
-In the MultiTask evaluation, each task represents a distinct skill domain (e.g., math, geoscience, mechanics). Let be the set of benchmark tasks.
+A.2.1 Application to ScienceEval Evaluation
+In the ScienceEval evaluation, each task represents a distinct skill domain (e.g., math, geoscience, biology). Let be the set of benchmark tasks.
 For each player and task , and for each round , we compute the percentage gain:
 We test each pair independently. To assess holistic advantage, we also aggregate over tasks:
-We consider RT statistically superior to player if:
+We consider RP statistically superior to player if:
 the Wilcoxon test yields on at least 8 of 9 tasks, or
 the aggregate test on satisfies .
-A.2.2 Application to SingleTask Evaluation
-In the SingleTask evaluation, each round produces rubric-based scores over 9 dimensions (covering three traits for each of Background, Methodology, and Impact).
+A.2.2 Application to ScienceNarrativeEvaluation
+In the ScienceNarrative evaluation, each round produces rubric-based scores over 9 dimensions (covering three traits for each of Background, Methodology, and Impact).
 For each player , round , and dimension , we compute:
@@
 Interpretation.
-The MultiTask analysis evaluates whether RT consistently improves performance across structurally diverse problems and domains.
-The SingleTask analysis inspects RT’s advantage across nuanced dimensions of academic writing quality.
-In both cases, aggregate-level tests offer a robust indicator of RT’s overall statistical superiority over individual LLM baselines.
+The ScienceEval analysis evaluates whether RP consistently improves performance across structurally diverse problems and domains.
+The ScienceNarrative analysis inspects RP’s advantage across nuanced dimensions of academic writing quality.
+In both cases, aggregate-level tests offer a robust indicator of RP’s overall statistical superiority over individual LLM baselines.
 A.3 Formal Framework for Kendall’s Tau Rank Correlation Analysis
@@
 Here indicates perfect agreement, perfect disagreement, and statistical independence of the rankings.
-Application to MultiTask.
-In the MultiTask setting, each task yields a separate ranking of player models by each grader.
+Application to ScienceEval.
+In the ScienceEval setting, each task yields a separate ranking of player models by each grader.
 For every round and task , we compute for all pairs of graders using their per-round rankings of the nine players.
@@
 where is the number of valid rounds for task . This produces a correlation matrix per task, summarizing grader consistency across domains such as mathematics, geoscience, and biology.
-Application to SingleTask.
-In the SingleTask setting, each round produces rubric-based scores across nine dimensions .
+Application to ScienceNarrative.
+In the ScienceNarrative setting, each round produces rubric-based scores across nine dimensions .
 For each dimension , graders generate rankings of the players, and we compute for all pairs .
@@
 Summary.
-In both MultiTask and SingleTask, Kendall’s Tau provides a principled measure of inter-grader consistency.
-The distinction lies in the evaluation unit: tasks for MultiTask versus rubric dimensions for SingleTask.
+In both settings, Kendall’s Tau provides a principled measure of inter-grader consistency.
+The distinction lies in the evaluation unit: tasks for ScienceEval versus rubric dimensions for ScienceNarrative.
 Appendix B Detailed Demonstration of Roundtable Policy
-To provide a concrete illustration of how Roundtable Policy (RT) operates in practice, we present two representative demonstrations: one in the MultiTask regime (Figure 7) and one in the SingleTask regime (Figure 6).
-These examples showcase how RT transforms fragmented outputs from multiple LLMs into coherent, factually grounded, and narratively rigorous responses.
-MultiTask demonstration.
-Figure 7 presents a representative MultiTask round, where nine LLM players respond simultaneously to a heterogeneous set of scientific queries spanning mechanics, thermodynamics, semiconductor physics, mathematics, geoscience, and biology.
+To provide a concrete illustration of how Roundtable Policy (RP) operates in practice, we present two representative demonstrations: one in the ScienceEval regime (Figure 9) and one in the ScienceNarrative regime (Figure 8).
+These examples showcase how RP transforms fragmented outputs from multiple LLMs into coherent, factually grounded, and narratively rigorous responses.
+ScienceEval demonstration.
+Figure 9 presents a representative ScienceEval round, where nine LLM players respond simultaneously to a heterogeneous set of scientific queries spanning mechanics, thermodynamics, semiconductor physics, mathematics, geoscience, and biology.
 Individual player outputs illustrate the limitations of single-model reasoning: some provide only partial explanations (e.g., noting that paper clips corrode without addressing stress relaxation), others introduce inaccuracies (e.g., oversimplifying nucleation–growth dynamics), and several fail to maintain consistency across domains (e.g., contradictory accounts of polymer glass transition).
 These fragmented outputs reflect the heterogeneity of LLM competencies—stronger players excel in certain areas, while weaker ones often produce vague or under-specified responses.
-By contrast, the fused response generated by RT demonstrates marked improvements in both factual coverage and reasoning depth.
-For instance, RT not only describes corrosion of paper clips but also contextualizes long-term degradation and mechanical implications; in the nucleation–growth problem, RT integrates complementary insights into thermodynamic driving forces and kinetic limitations, producing a textbook-quality explanation.
-Similarly, in the polymer glass transition question, RT combines molecular structure, chain length, and plasticizer effects into a unified account that no single player captured in full.
-Committee grading further substantiates these gains: RT exhibits higher factual accuracy and reduced uncertainty intervals across tasks.
-Improvements are particularly pronounced in symbolic and reasoning-intensive domains such as thermodynamics and mathematics, where RT delivers correct solutions while clarifying intermediate reasoning steps.
-These results highlight how RT balances correctness and clarity across heterogeneous domains.
-SingleTask demonstration.
-Figure 6 illustrates the SingleTask regime, where each player generates a three-paragraph scientific proposal (Background, Methodology, Impact).
+By contrast, the fused response generated by RP demonstrates marked improvements in both factual coverage and reasoning depth.
+For instance, RP not only describes corrosion of paper clips but also contextualizes long-term degradation and mechanical implications; in the nucleation–growth problem, RP integrates complementary insights into thermodynamic driving forces and kinetic limitations, producing a textbook-quality explanation.
+Similarly, in the polymer glass transition question, RP combines molecular structure, chain length, and plasticizer effects into a unified account that no single player captured in full.
+Committee grading further substantiates these gains: RP exhibits higher factual accuracy and reduced uncertainty intervals across tasks.
+Improvements are particularly pronounced in symbolic and reasoning-intensive domains such as thermodynamics and mathematics, where RP delivers correct solutions while clarifying intermediate reasoning steps.
+These results highlight how RP balances correctness and clarity across heterogeneous domains.
+ScienceNarrative demonstration.
+Figure 8 illustrates the ScienceNarrative regime, where each player generates a three-paragraph scientific proposal (Background, Methodology, Impact).
 Individual drafts vary widely in creativity, methodological depth, and logical flow.
@@
 Such stylistic and structural inconsistencies exemplify the difficulty of producing a coherent scientific narrative with a single LLM.
-In contrast, the fused proposal produced by RT exhibits a markedly stronger narrative arc.
+In contrast, the fused proposal produced by RP exhibits a markedly stronger narrative arc.
 The Background situates the research problem within its broader scientific context; the Methodology clearly articulates technical contributions with precision; and the Impact section highlights both scientific significance and societal relevance.
-Grading results confirm consistent improvements across all rubric dimensions—creativity, scientific rigor, and logical coherence—demonstrating that RT effectively mitigates individual stylistic biases while amplifying collective strengths.
-Even strong baseline players benefit: while their drafts are often polished in one dimension, RT integrates complementary insights to deliver uniformly high-quality writing.
+Grading results confirm consistent improvements across all rubric dimensions—creativity, scientific rigor, and logical coherence—demonstrating that RP effectively mitigates individual stylistic biases while amplifying collective strengths.
+Even strong baseline players benefit: while their drafts are often polished in one dimension, RP integrates complementary insights to deliver uniformly high-quality writing.
 Summary.
 Together, these demonstrations show how Roundtable Policy transforms fragmented or uneven model outputs into coherent, high-quality responses.
-In the MultiTask regime, RT enhances cross-domain reasoning reliability by integrating complementary knowledge across scientific fields.
-In the SingleTask regime, RT substantially improves narrative structure and explanatory depth in long-form proposal writing.
-These examples reinforce the central claim that RT enables LLM ensembles to achieve more factual, rigorous, and narratively coherent outputs in scientific applications.
+In the ScienceEval regime, RP enhances cross-domain reasoning reliability by integrating complementary knowledge across scientific fields.
+In the ScienceNarrative regime, RP substantially improves narrative structure and explanatory depth in long-form proposal writing.
+These examples reinforce the central claim that RP enables LLM ensembles to achieve more factual, rigorous, and narratively coherent outputs in scientific applications.
 Appendix C Benchmarking Datasets
 C.1 Dataset Annotation
-We introduce MatSci–Bench, a heterogeneous benchmark designed to evaluate the scientific reasoning capabilities of large language models (LLMs) in materials science and beyond.
+We introduce Multi-Faceted Physical Reasoning, a heterogeneous benchmark designed to evaluate the scientific reasoning capabilities of LLMs in physical science and beyond.
 The benchmark spans six carefully curated sub-domains: (i) Macro Mechanics, (ii) Micro Mechanics, (iii) Optical & Microscopic, (iv) Semiconductor Physics, (v) Thermodynamical Systems, and (vi) Electrical & Electronic Properties.
@@
 These fine-grained metrics allow for robust measurement of model performance and principled ablation studies.
-Table 4 provides representative examples across four broader scientific areas—Biology, Geoscience, Mathematics, and Materials Science—highlighting domain-specific instruction design and diverse task objectives.
+Table 5 provides representative examples across four broader scientific areas—Biology, Geoscience, Mathematics, and Physical Science—highlighting domain-specific instruction design and diverse task objectives.
 Unlike prior benchmarks focused narrowly on short factual queries, our dataset emphasizes heterogeneous scientific tasks requiring numerical derivation, mechanistic explanation, and symbolic fluency.
-Materials-science subset.
-The materials-science questions are drawn from the 2006–2024 UCLA Materials Science and Engineering Ph.D.
+Multi-Faceted Physical Reasoning.
+The physical-science questions are drawn from the 2006–2024 UCLA Ph.D.
 qualifying examinations.
@@
 All annotations are fully compliant with LaTeX and Markdown standards, ensuring reproducibility and integration with downstream evaluation pipelines.
-Subdomain Task Instruction Subtask Math Task 007: Math Janet’s ducks lay 16 eggs per day.
+Subdomain Task Instruction More details Math Task 007: Math Janet’s ducks lay 16 eggs per day.
 She consumes 3 eggs for breakfast, uses 4 eggs for baking muffins, and sells the remaining eggs at $2 each.
@@
 Link pharmacological interventions with clinical outcomes based on empirical data.
-Materials Science Macro Mechanics Task 001: Mech.
+Physical Science Macro Mechanics Task 001: Mech.
 1 What is elasticity?
@@
 Demonstrate a solid understanding of quantum mechanics by providing the correct form of Schrödinger’s equation and a step-by-step solution for a 1D potential problem.
-Success Criteria: The arbitrator agent confirms that the confidence score of the multi-agents converges.
 C.2 Dataset Comparison
 | Domain | Q/K | Tokens |
-| Geoscience | 1395 | 109,159 |
-| Biology | 1000 | 92,866 |
-| Math | 1319 | 246,238 |
-| Material Science | 539 | 46,150 |
-Table 5 benchmarks MatSci–Bench against a range of prominent scientific QA datasets.
-Whereas prior benchmarks often emphasize scale—providing thousands of short, loosely categorized questions—MatSci–Bench is designed for conceptual depth, compositional reasoning, and graduate-level rigor.
-For instance, the materials science subset averages tokens per question, more than double that of typical mathematics datasets, reflecting the structural and semantic complexity encountered in authentic scientific workflows.
+| Geoscience(Deng et al., 2024) | 1395 | 109,159 |
+| Biology(Jin et al., 2019) | 1000 | 92,866 |
+| Math(Cobbe et al., 2021) | 1319 | 246,238 |
+| Physical Science | 539 | 46,150 |
+Table 6 benchmarks Multi-Faceted Physical Reasoning against a range of prominent scientific QA datasets.
+Whereas prior benchmarks often emphasize scale—providing thousands of short, loosely categorized questions—Multi-Faceted Physical Reasoning is designed for conceptual depth, compositional reasoning, and graduate-level rigor.
+For instance, the physical science subset averages tokens per question, more than double that of typical mathematics datasets, reflecting the structural and semantic complexity encountered in authentic scientific workflows.
 To construct this benchmark, we developed a hybrid annotation pipeline that balances automated scalability with human quality control.
@@
 We evaluate the framework under two complementary protocols:
-MultiTask.
+ScienceEval.
 Each round presents a batch of heterogeneous tasks spanning multiple scientific domains as shown in Figure 11.
 The AI Committee jointly evaluates player responses, testing the arbitration system’s ability to generalize across diverse input types, modalities, and difficulty levels.
-SingleTask.
+ScienceNarrative.
 Each round focuses on a scientific proposal generation as shown in Figure 12, allowing the Committee to learn nuanced arbitration strategies within a fixed domain and structured output format.
-Together, these protocols provide a comprehensive evaluation of arbitration dynamics under both breadth-oriented (MultiTask) and depth-oriented (SingleTask) conditions.
+Together, these protocols provide a comprehensive evaluation of arbitration dynamics under both breadth-oriented (ScienceEval) and depth-oriented (ScienceNarrative) conditions.
 Importantly, all evaluations are conducted without modifying the internal parameters of the player models, thereby preserving their pretrained competencies while enabling compositional reasoning through learned arbitration.
-C.4 MatSci–Bench
-Table 6 illustrates the mapping between several UCLA courses (from the prelim exams spanning 2006–2024) and our six defined subdomains in Material Science.
+C.4 Multi-Faceted Physical Reasoning
+Table 7 illustrates the mapping between several UCLA courses (from the prelim exams spanning 2006–2024) and our six defined subdomains in Multi-Faceted Physical Reasoning.
 The table indicates the number of questions per course and marks (✓) the corresponding subdomains covered by each course.
-And Table 7 Mapping from numeric course codes to official course names in the UCLA MSE Graduate Program.
-In our multi-task framework, each question is paired with a corresponding key—forming a question-key pair—that encapsulates its unique characteristics and domain-specific attributes.
+And Table 8 Mapping from numeric course codes to official course names.
+In ScienceEval, each question is paired with a corresponding key—forming a question-key pair—that encapsulates its unique characteristics and domain-specific attributes.
 This design enables a fine-grained classification of tasks, as each key serves as an identifier for the subdomain and the underlying problem-solving facet it represents.
-In our dataset, which is constructed from high-quality, authoritative UCLA prelim exams, each question is not only rigorously structured but also assigned a key that maps to distinct subdomains within Material Science.
-This approach allows for a combinatorial evaluation across various tasks and provides a nuanced understanding of the performance of large language models (LLMs).
+In our dataset, which is constructed from high-quality, authoritative UCLA prelim exams, each question is not only rigorously structured but also assigned a key that maps to distinct subdomains within Physical Science.
+This approach allows for a combinatorial evaluation across various tasks and provides a nuanced understanding of the performance of LLMs.
 Moreover, the explicit association between questions and keys facilitates a higher resolution in assessment, offering insights into the model’s capability to address the subtle distinctions inherent in each category.
@@
 Appendix D Fine-grained Results
-D.1 MultiTask
+D.1 ScienceEval
 Player ID Task ID # (%) (%) p-value Significance Overall Player_001 GPT-4o 001 36000 3.664534 29.115645 ✓ ✓ 002 36000 9.990613 66.756230 ✓ 003 36000 4.418581 32.628933 ✓ 004 36000 18.290543 103.900081 ✓ 005 36000 3.507243 49.094848 ✓ 006 36000 4.547669 29.876047 ✓ 007 36000 18.619909 110.561195 ✓ 008 36000 4.205786 76.940468 ✓ 009 36000 -13.817889 121.279772 ✗ Player_002 Grok-3 001 36000 2.442676 26.869884 ✓ ✓ 002 36000 11.658807 87.033589 ✓ 003 36000 2.922847 25.540567 ✓ 004 36000 10.040225 97.107301 ✓ 005 36000 2.754232 68.334607 ✓ 006 36000 7.795831 50.756077 ✓ 007 36000 7.037656 73.901148 ✓ 008 36000 12.159710 90.508914 ✓ 009 36000 -14.196161 124.567204 ✗ Player_003 Claude-3.5Haiku 001 36000 9.356006 49.469078 ✓ ✓ 002 36000 25.634084 100.488313 ✓ 003 36000 8.903574 52.492393 ✓ 004 36000 24.974510 120.272436 ✓ 005 36000 18.091843 89.503620 ✓ 006 36000 19.206568 71.695176 ✓ 007 36000 29.644628 127.261225 ✓ 008 36000 22.433517 119.787354 ✓ 009 36000 -8.057303 138.391228 ✗ Player_004 GPT-4.5-preview 001 36000 26.864778 50.638001 ✓ ✓ 002 36000 24.364182 76.170729 ✓ 003 36000 28.888756 93.939402 ✓ 004 36000 18.305048 104.942899 ✓ 005 36000 18.611880 82.101495 ✓ 006 36000 17.971459 45.269719 ✓ 007 36000 32.007776 136.936369 ✓ 008 36000 5.146633 83.614309 ✓ 009 36000 -11.464038 128.061246 ✗ Player_005 Deepseek-chat 001 36000 15.448390 50.327417 ✓ ✓ 002 36000 21.685092 88.100774 ✓ 003 36000 11.772803 42.894447 ✓ 004 36000 14.794087 99.773548 ✓ 005 36000 9.123905 59.031496 ✓ 006 36000 12.303097 45.278523 ✓ 007 36000 52.148935 167.006012 ✓ 008 36000 11.541473 99.093907 ✓ 009 36000 -8.504933 137.294744 ✗
 Player ID Task ID # (%) (%) p-value Significance Overall Player_006 Claude-3.7Sonnet 001 36000 10.483677 100.447956 ✓ ✓ 002 36000 20.034584 140.497283 ✓ 003 36000 7.986260 57.805017 ✓ 004 36000 11.881086 139.264504 ✗ 005 36000 4.431706 106.492570 ✓ 006 36000 11.670759 84.754451 ✓ 007 36000 65.488159 182.402404 ✓ 008 36000 10.358145 91.889615 ✓ 009 36000 -16.646343 131.246587 ✗ Player_007 Llama-3.3 001 36000 26.016269 81.867305 ✓ ✓ 002 36000 43.795534 133.423412 ✓ 003 36000 29.847791 118.361082 ✓ 004 36000 33.835778 146.064900 ✓ 005 36000 31.906208 125.420480 ✓ 006 36000 29.654251 95.808593 ✓ 007 36000 12.560069 113.533051 ✓ 008 36000 24.151429 138.741394 ✓ 009 36000 -8.214252 139.363305 ✗ Player_008 Qwen-plus 001 36000 12.702783 53.480581 ✓ ✓ 002 36000 24.671245 105.447534 ✓ 003 36000 12.217150 54.159535 ✓ 004 36000 19.478679 117.267431 ✓ 005 36000 6.452739 58.406561 ✓ 006 36000 18.500964 66.688703 ✓ 007 36000 16.654888 101.522057 ✓ 008 36000 15.475509 98.223206 ✓ 009 36000 -13.418640 129.270826 ✗ Player_009 Gemini-2.0 001 36000 48.597322 101.645241 ✓ ✓ 002 36000 56.207001 136.978257 ✓ 003 36000 30.807297 86.405085 ✓ 004 36000 40.730016 138.029856 ✓ 005 36000 39.623737 120.151633 ✓ 006 36000 43.606884 96.069039 ✓ 007 36000 127.050040 253.842133 ✓ 008 36000 19.718319 112.418143 ✓ 009 36000 15.955998 173.200269 ✓
-D.2 SingleTask
+D.2 ScienceNarrative
 Player ID Paragraph Dimension # (%) (%) p-value Significance Overall Player_001 GPT-4o Background Creativity 36000 37.150702 49.336875 ✓ ✓ Scientific Rigor 36000 30.257828 46.850598 ✓ Logical Coherence 36000 11.733576 16.004795 ✓ Methodology Creativity 36000 24.931209 31.921273 ✓ Scientific Rigor 36000 31.838154 45.990879 ✓ Logical Coherence 36000 17.282977 21.020188 ✓ Impact Creativity 36000 32.606157 40.063019 ✓ Scientific Rigor 36000 41.271411 57.887977 ✓ Logical Coherence 36000 16.718015 19.804013 ✓ Player_002 GPT-o3 Background Creativity 36000 12.164463 112.053304 ✓ ✓ Scientific Rigor 36000 13.556005 141.849680 ✓ Logical Coherence 36000 7.902093 88.022126 ✓ Methodology Creativity 36000 12.242360 162.582687 ✓ Scientific Rigor 36000 13.742029 176.441027 ✓ Logical Coherence 36000 12.246663 149.530625 ✓ Impact Creativity 36000 14.822979 162.636940 ✓ Scientific Rigor 36000 19.794367 172.889406 ✓ Logical Coherence 36000 13.375266 149.888340 ✓ Player_003 GPT-4.1 Background Creativity 36000 23.490292 76.488563 ✓ ✓ Scientific Rigor 36000 15.516892 73.369125 ✓ Logical Coherence 36000 7.636895 70.613455 ✓ Methodology Creativity 36000 13.841302 73.356046 ✓ Scientific Rigor 36000 13.786087 74.585712 ✓ Logical Coherence 36000 9.985966 70.810643 ✓ Impact Creativity 36000 18.657843 73.142286 ✓ Scientific Rigor 36000 22.049522 75.719467 ✓ Logical Coherence 36000 11.006198 72.520043 ✓ Player_004 GPT-4.1-nano Background Creativity 36000 22.875212 41.517104 ✓ ✓ Scientific Rigor 36000 16.303948 31.393619 ✓ Logical Coherence 36000 7.715497 14.464709 ✓ Methodology Creativity 36000 14.951288 21.488109 ✓ Scientific Rigor 36000 14.777490 25.696003 ✓ Logical Coherence 36000 10.347582 16.738338 ✓ Impact Creativity 36000 19.842341 27.921968 ✓ Scientific Rigor 36000 23.655715 38.425349 ✓ Logical Coherence 36000 11.784470 18.251267 ✓ Player_005 GPT-o3-mini Background Creativity 36000 12.773348 27.212152 ✓ ✓ Scientific Rigor 36000 9.060914 24.819729 ✓ Logical Coherence 36000 5.406290 14.736602 ✓ Methodology Creativity 36000 10.166683 16.487638 ✓ Scientific Rigor 36000 7.171348 19.411696 ✓ Logical Coherence 36000 6.591953 13.835775 ✓ Impact Creativity 36000 13.300165 22.726699 ✓ Scientific Rigor 36000 14.308252 31.344613 ✓ Logical Coherence 36000 8.408882 16.386089 ✓
@@
 For example, a grader may consistently assign lower (more critical) ranks, while another may tend to be more lenient.
-If such biases are not controlled, downstream evaluations may unfairly penalize or reward certain systems, obscuring the true effect of the Roundtable Policy.
+If such biases are not controlled, downstream evaluations may unfairly penalize or reward certain systems, obscuring the true effect of Roundtable Policy.
 Consistency, on the other hand, reflects whether different graders produce comparable relative orderings.
@@
 We assess grader bias by visualizing the empirical rank distributions assigned by each arbitrator.
-Figure 3 presents kernel density estimates of the rank distributions for four graders across both the SingleTask and MultiTask evaluations.
+Figure 7 presents kernel density estimates of the rank distributions for four graders across both the ScienceEval and ScienceNarrative evaluations.
 Systematic deviations in the distributions (e.g., skewed toward higher or lower ranks) indicate bias.
@@
 Findings.
-In the SingleTask setting, we observe substantial variation in grader behavior.
+In the ScienceNarrative, we observe substantial variation in grader behavior.
 The rank distributions are uneven, with some graders systematically harsher or more generous than others.
@@
 Such findings echo common phenomena in human peer review, where individual reviewers bring idiosyncratic standards and biases.
-By contrast, in the MultiTask setting, the distributions are more balanced, and the Kendall’s Tau values are consistently higher.
+By contrast, in the ScienceEval, the distributions are more balanced, and the Kendall’s Tau values are consistently higher.
 This suggests that when grading across diverse tasks, individual biases are averaged out, leading to more stable and reliable inter-grader consensus.
@@
 First, reliance on a single grader is inherently risky due to potential bias; employing multiple graders and aggregating their judgments is essential for fairness.
-Second, broader task coverage increases reliability, as MultiTask grading mitigates the influence of any single arbitrator’s bias.
+Second, broader task coverage increases reliability, as ScienceEval grading mitigates the influence of any single arbitrator’s bias.
 Together, these results highlight the necessity of designing evaluation protocols that balance scalability with robustness, ensuring that AI grader-based assessments approximate the fairness and credibility traditionally associated with human expert panels.
+Appendix E Multi-Agent Baselines
+Majority Vote (Wang et al., 2023).
+In the majority vote baseline, multiple agents independently generate candidate responses to the same query.
+The final output is selected by simple majority voting over the agents’ responses.
+No historical information or agent-specific weighting is used; all agents contribute equally at inference time.
+Weighted Vote (Taubenfeld et al., 2025).
+Weighted vote extends majority voting by assigning different weights during aggregation.
+In our implementation, weights are determined by self-reported confidence scores provided by the agents.
+The final prediction is obtained through confidence-weighted aggregation rather than uniform voting.
+Debate (Du et al., 2024).
+Debate is an interaction-based multi-agent framework in which agents iteratively exchange arguments to refine their responses.
+Agents begin by independently generating initial answers, followed by multiple rounds of critique and revision.
+After the final debate round (three rounds in our implementation), the system outputs one agent’s response, corresponding to the final updated solution.
+Debate w Judge (Liang et al., 2024).
+Debate w Judge augments the standard debate framework by introducing an explicit adjudicator agent.
+Following several rounds of debate (three rounds in our implementation), the adjudicator reviews all final responses and selects the most convincing answer.
+To ensure architectural consistency across methods, we employ the same aggregator model used in Roundtable Policy as the judge.
+The adjudicator is applied only at the final decision stage and does not intervene during the debate process itself.
+Debate w Confidence (Chen et al., 2024).
+Debate w Confidence further extends debate by incorporating agent confidence signals into the aggregation process.
+The framework proceeds in three phases.
+First, each agent independently generates an initial response and accompanying explanation.
+Second, agents engage in multiple rounds of discussion (three rounds in our implementation), updating their responses based on information exchanged with peers.
+As both the ScienceEval and ScienceNarrative settings involve open-ended reasoning rather than discrete choices, the grouping mechanism is instantiated as concatenation of peer responses.
+Finally, after each discussion round, the system produces an output via confidence-weighted voting over agent responses.
+Mixture-of-Agents (Wang et al., 2025).
+Mixture-of-Agents is a layered multi-agent aggregation framework that exploits the observed collaborativeness of large language models, whereby response quality can improve when conditioned on outputs from other agents.
+In our implementation, we adopt a three-layer architecture with nine agents per layer.
+The same aggregator model as in Roundtable Policy is used for the final aggregation stage to maintain consistency across methods.